### v1

# Berry phases and connection matrices defined on homogeneous spaces attached to Siegel-Jacobi groups

Abstract.
The relation between the Berry phase and connection matrix on the Siegel-Jacobi disk ${\mathcal{D}}^{J}_{1}$ and Siegel-Jacobi upper half-plane ${\mathcal{X}}^{J}_{1}$ are analyzed. The connection matrix and the covariant derivative of one-forms on the extended Siegel-Jacobi upper half-plane $\tilde{{\mathcal{X}}}^{J}_{1}$ are calculated.
Key words and phrases:
Berry phase, Berry connection, Connection matrix, Jacobi group, invariant metric, Siegel–Jacobi disk, Siegel–Jacobi upper half-plane, extended Siegel–Jacobi upper half-plane, almost cosymplectic manifold
1991 Mathematics Subject Classification:
81Q70,53C05,37J55,53D15
1. Introduction
The complex Jacobi group [56, 40] of index $n$ is defined as the semi-direct product $G^{J}_{n}={\mathrm{H}}_{n}\rtimes\text{Sp}(n,\mathbb{R})_{\mathbb{C}}$, where ${\mathrm{H}}_{n}$ denotes the $(2n+1)$-dimensional Heisenberg group [90, 18, 19]. To the Jacobi group $G^{J}_{n}$ it is associated a homogeneous manifold, called the Siegel-Jacobi ball ${\mathcal{D}}^{J}_{n}$ [18], whose points are in $\mathbb{C}^{n}\times{\mathcal{D}}_{n}$, i.e. a partially-bounded space [93, 94]. ${\mathcal{D}}_{n}$ denotes the Siegel (open) ball of index $n$. The non-compact Hermitian symmetric space $\operatorname{Sp}(n,\mathbb{R})_{\mathbb{C}}/\operatorname{U}(n)$ admits a matrix realization as a homogeneous bounded domain [60]:
${\mathcal{D}}_{n}:=\{W\in MS(n,\mathbb{C}):\mathbb{1}_{n}-W\bar{W}>0\}.$
The real Jacobi group of degree $n$ is defined as $G^{J}_{n}(\mathbb{R}):={\rm Sp}(n,\mathbb{R})\ltimes{\mathrm{H}}_{n}$, where ${\mathrm{H}}_{n}={\mathrm{H}}_{n}(\mathbb{R})$ is the real $(2n+1)$-dimensional Heisenberg group. ${\rm Sp}(n,\mathbb{R})_{\mathbb{C}}$ and $G^{J}_{n}$ are isomorphic to ${\rm Sp}(n,\mathbb{R})$ and $G^{J}_{n}(\mathbb{R})$ respectively as real Lie groups, see [19, Proposition 2].
The invariant metric on the Siegel-Jacobi upper half-plane on ${\mathcal{X}}^{J}_{1}=\frac{G^{J}_{1}(\mathbb{R})}{{\rm SO}(2)\times\mathbb{R}}\approx{\mathcal{X}}_{1}\times\mathbb{R}^{2}$ [15, 16, 21, 22] was obtained previously by Berndt [39, 38] and Kähler [61, 62].
We determined the invariant metric on a five dimensional homogeneous manifold ${\tilde{{\mathcal{X}}}^{J}_{1}}=\frac{G^{J}_{1}(\mathbb{R})}{{\rm SO}(2)}\approx{\mathcal{X}}_{1}\times\mathbb{R}^{3}$ [27], called the extended Siegel–Jacobi upper half-plane. The results of [27] concerning $\tilde{{\mathcal{X}}}^{J}_{1}$ have been generalized in [28] to the extended Siegel-Jacobi upper half space $\tilde{{\mathcal{X}}}^{J}_{n}=\frac{G^{J}_{n}(\mathbb{R})}{\text{U}(n)}\approx{\mathcal{X}}^{J}_{n}\times\mathbb{R}$, ${\mathcal{X}}^{J}_{n}\approx\mathbb{C}^{n}\times{\mathcal{X}}_{n}$, ${\mathcal{X}}_{n}=\frac{{\mathrm{Sp}}(n,\mathbb{R})}{{\mathrm{U}}(n)},$ $\mathbb{N}\ni n>1$.
We recall that on homogenous Kähler  manifolds the Hamilton equations of motion and the Berry phase were simultaneously investigated in [35, 31, 21], see also [57]. In the present paper we are interested in the same problem of studying the Berry phase on odd-dimensional manifolds, where several geometric structures can be introduced [41, 42, 43, 44, 46, 47, 67, 69, 82], see also a brief review in [30, Appendix]. In our paper [30] we have investigated Hamiltonian systems on manifolds with almost cosymplectic structure in the sense of [70]. In the present paper we investigate the connection matrix on odd dimensional manifolds endowed with an almost complex structure.
We recall here our interrest to find a geometric significance to the phase of the scalar product of coherent states [79, 71, 75]. The answer to this question was given by Pancharatnam for the Poincaré sphere [78, 84], see also [1] and [74, Proposition 5.1] in the language of holonomy (see § 5.1.3) of a loop in the projective Hilbert space, and by Perelomov [79, page 63] for the sphere $\text{S}^{2}=\frac{\text{SU}(2)}{\text{U}(1)}$. A general answer to this question using the coherent state embedding and the so called ” Cauchy formulas” was given in [11] and [37]. We also studied this problem in [9][12]. Explicit calculation was presented for the compact Grassmann manifold $G_{n}(\mathbb{C}^{m+n})=\frac{\text{SU}(n+m)}{\text{S}(\text{U}(n)\times\text{U}(m))}$ in [11], where it was proved that the phase of the scalar product of two coherent states is twice the symplectic area of a geodesic triangle determined by the corresponding points on the manifold and the origin of the system of coordinates, see also [51, Theorem 2.1]. The same result is also true for the noncompact dual $\frac{\text{SU}(n,m)}{\text{S}(\text{U}(n)\times\text{U}(m))}$ of the compact Grassmann manifold [9, 10]. In [21] the change of coordinates $x\rightarrow z$ in (3.1) below was called $FC$-transform (fundamental conjecture [86, 52, 55]). We observed that [31, Remark 3]
For symmetric manifolds the FC-transform gives geodesics  (A)
In [8, Remark 1] we underlined that assertion (A) is true for class of manifolds which includes the naturally reductive spaces [76, 3], [64, page 202]. We have considered the sequence of manifolds
Hermitian symmetric spaces $\subset$ symmetric $\subset$ naturally reductive $\subset$ g. o.
We have shown in [27, Proposition 5.8] that ${\mathcal{X}}^{J}_{1}$ is not naturally reductive with respect to the balanced metric [54, 2, 72]. In [29, Theorem 1] we have proved that the extended Siegel-Jacobi upper half-plane, realized as homogenous Riemannian manifold ($\tilde{{\mathcal{X}}}^{J}_{1}=\frac{G^{J}_{1}(\mathbb{R})}{\text{SO}(2)},g_{\tilde{{\mathcal{X}}}^{J}_{1}}$) is a reductive, non-symmetric manifold, non-naturally reductive with respect with the metric (2.51), not a g.o. space [66] with respect to the invariant metric $g_{\tilde{{\mathcal{X}}}^{J}_{1}}$.
We recall that the Berry phase is an important object in the study of geometric phase physics [83, 84, 57, 77]. We have studied the Berry phase on homogenous Kähler  manifolds in [35, 31, 21].
The paper is laid out as follows. In Section 2 we recall the Kähler  two-form on ${\mathcal{D}}^{J}_{1}$ and its two-parameter balanced metric image on ${\mathcal{X}}^{J}_{1}$ obtained by the partial Cayley transform in Proposition 1, the three parameter invariant metric on $\tilde{{\mathcal{X}}}^{J}_{1}$ in the S-coordinates [83] in Proposition 2, while Proposition 3 recalls the invariant metric on ${\mathcal{D}}^{J}_{n}$, ${\mathcal{X}}^{J}_{n}$ and $\tilde{{\mathcal{X}}}^{J}_{n}$. Section 3 recalls our investigation on Berry phase on Kähler  manifolds. In particular, Proposition 4 recalls the Berry phase on ${\mathcal{D}}^{J}_{1}$ and ${\mathcal{X}}^{J}_{1}$. Section 4 summarise the notion of almost cosymplectic manifold [70]. In particular, the manifold $\tilde{{\mathcal{X}}}^{J}_{1}$ is endowed with a generalized transitive almost cosymplectic structure [30]. In §4.2 is calculated the connection matrix [53] on $\tilde{{\mathcal{X}}}^{J}_{1}$ and in §4.3 are presented the covariant derivative (as in [88, 89], see also [2, §3.2]) on ${{\mathcal{X}}}^{J}_{1}$ and $\tilde{{\mathcal{Xs}}}^{J}_{1}$. The last Section - Appendix - collects several mathematical concepts used in the paper: connections on real manifolds in Subsection 5.1.1, connections on complex manifolds, Chern connections [48] and quantizable Kähler  manifolds in Subsection 5.1.2, the notion of holonomy [65] is recalled in Subsection 5.1.3. Some example are contained in Subsection 5.2 devoted to coherent states: Berry connection and Kähler  two-form for the Heisenberg-Weyl group, sphere $S^{2}$, ${\mathcal{D}}_{1}$, complex Grassmann manifold $G_{n}(\mathbb{C}^{n+m})$ and its non-compact dual, ${\mathbb{CP}}^{n}$ and ${\mathbb{CP}}^{n,1}$.
Proposition 1 and Comment 1 are improved versions of older results, while Remark 3 compare our approach to Berry phase on Kähler  manifolds [35, 31, 21] with the geometric phase in [84, 83, 50]. The new relevant results presented in this paper are § 3.4, formula (3.47) of $\theta_{{\mathcal{X}}^{J}_{1}}(x,y,q,p)$, formula (3.66) of Berry phase on ${\mathcal{X}}^{J}_{1}$ in $(u,v)=(m+\operatorname{i}n,x+\operatorname{i}y)$, formulae (3.67), (3.7), (3.7) of the Berry phase in $(w,v)=(\alpha+\operatorname{i}\beta,x+\operatorname{i}y)$, (3.70) for the Berry phase in $(x,y,q,p)$, Lemma 2 which gives $\omega_{{\mathcal{X}}^{J}_{n}}(x,y,p,q)$, formula (4.2) $\theta^{\prime}_{\tilde{{\mathcal{X}}}^{J}_{1}}(x,y,q,p,\kappa)$ of the connection matrix on $\tilde{{\mathcal{X}}^{J}_{1}}$, the covariant derivatives $Dx$,( $Dy,~{}Dq,~{}Dp$) (4.25), ( (4.26), (4.27), respectively (4.28)) on ${\mathcal{X}}^{J}_{1}$, formulae of $Dx$, ($Dy,~{}Dq,~{}Dp,~{}D\kappa$) (4.29) ((4.30), (4.31), (4.32), respectively (4.33)) on $\tilde{{\mathcal{X}}}^{J}_{1}$.
Notation We denote by $\mathbb{R}$, $\mathbb{C}$, $\mathbb{Z}$ and $\mathbb{N}$ the field of real numbers, the field of complex numbers, the ring of integers, and the set of non-negative integers, respectively. We denote the imaginary unit $\sqrt{-1}$ by $\operatorname{i}$, the real and imaginary parts of a complex number $z\in\mathbb{C}$ by $\operatorname{Re}z$ and $\operatorname{Im}z$ respectively, and the complex conjugate of $z$ by $\bar{z}$. We denote by ${\operatorname{d}}$ the differential. We use Einstein’s summation convention, i.e. repeated indices are implicitly summed over. The set of vector fields (1-forms) on real manifolds is denoted by ${\mathfrak{D}}^{1}$ (respectively ${\mathfrak{D}}_{1}$). We denote a mixed tensor contravariant of degree $r$ and covariant of degree $s$ by ${\mathfrak{D}}^{r}_{s}={\mathfrak{D}}^{r}\times{\mathfrak{D}}_{s}$, where ${\mathfrak{D}}^{r}=\underbrace{{\mathfrak{D}}^{1}\times\dots\times{\mathfrak{D}}^{1}}_{r}$ and ${\mathfrak{D}}_{s}=\underbrace{{\mathfrak{D}}_{1}\times\dots\times{\mathfrak{D}}_{1}}_{s}$ [60, pages 13-17]. If $M$ is a complex manifold we denote by ${\mathfrak{A}}^{r,s}$ the tensor fields of type $(r,s)$. If we denote with Roman capital letteres the Lie groups, then their associated Lie algebras are denoted with the corresponding lower-case letteres. If ${\mathfrak{H}}$ is a Hilbert space, than we adopt the convention that the scalar product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on ${\mathfrak{H}}\times{\mathfrak{H}}$ is antilinear in the first factor $(\lambda a,b)=\bar{\lambda}(a,b),\quad\lambda\in\mathbb{C}\setminus\{0\}$. If $\pi$ is a representation of a Lie grup $G$ on the Hilbert ${\mathfrak{H}}$ and $X\in{\mathfrak{g}}$, then we denote $\bf{X}:=\operatorname{d}\pi(X)$ [14, 22, 79]. The interior product $i_{X}\omega$ (interior multiplication or contraction) of the differential form $\omega$ with $X\in{\mathfrak{D}}^{1}$ is denoted $X\lrcorner\omega$. We denote by $M(n,m,{\mathbb{F}})$ the set of $n\times m$ matrices with elements in the field ${\mathbb{F}}$ and $M(n,{\mathbb{F}})$ denotes $M(n,n,{\mathbb{F}})$. If $X\in M(n,m,{\mathbb{F}})$, then $X^{t}$ denotes the transpose of $X$. We denote by $MS(n,{\mathbb{F}})=\{X\in M(n,{\mathbb{F}})|X=X^{t}\}$. The conjugate transpose (or hermitian transpose) of $A\in M(q,\mathbb{C})$ is $A^{H}:=\bar{A}^{t}$, also denoted $A^{*}$, $A^{\dagger}$, $A^{+}$. If $f$ is a function on $\mathbb{C}^{n}$, we write for the total differential of $f$ $\operatorname{d}f=\partial f+\bar{\partial}f$, $\partial f=\sum_{1}^{n}{\partial_{\alpha}f}\operatorname{d}z_{\alpha}$, where $\partial_{\alpha}f=\frac{\partial f}{\partial z_{\alpha}}$ [59, page 6]. If $f$ is a complex function, then by $f-{cc}$ we mean $f-\bar{f}$.
2. Preparation
We adopt the notation from [40, 56] for the real Jacobi group $G^{J}_{1}(\mathbb{R})$, realized as submatrices of $\text{Sp}(2,\mathbb{R})$ of the form
(2.1) $g=\left(\begin{array}[]{cccc}a&0&b&q\\ \lambda&1&\mu&\kappa\\ c&0&d&-p\\ 0&0&0&1\end{array}\right),~{}M=\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right),~{}\det M=1,$
where
(2.2) $Y:=(p,q)=XM^{-1}=(\lambda,\mu)\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)^{-1}=(\lambda d-\mu c,-\lambda b+\mu a)$
is related to the Heisenberg group $\rm{H}_{1}$ described by $(\lambda,\mu,\kappa)$. For coordinatization of the real Jacobi group we adopt the so called $S$-coordinates $(x,y,\theta,p,q,\kappa)$ [40].
Simultaneously with the Jacobi group $G^{J}_{1}(\mathbb{R})$ consisting of elements $(M,X,\kappa)$, we considered the restricted real Jacobi group $G^{J}(\mathbb{R})_{0}$ of elements $(M,X)$ [15, 27].
The action $G^{J}(\mathbb{R})_{0}\times{\mathcal{X}}^{J}_{1}\rightarrow{\mathcal{X}}^{J}_{1}$ (respectively $G^{J}_{1}(\mathbb{R})\times\tilde{{\mathcal{X}}}^{J}_{1}\rightarrow\tilde{{\mathcal{X}}}^{J}_{1}$) in Lemma 1 below is extracted from [27, Lemma 5.1], [4, Lemma 1].
Let
(2.3) $\mathbb{C}\ni v:=x+\operatorname{i}y,~{}~{}~{}\mathbb{C}\ni u:=pv+q=\xi+\operatorname{i}\rho,\quad~{}x,y,p,q,\xi,\rho\in\mathbb{R}.$
Kähler calls $\tilde{{\mathcal{X}}}^{J}_{1}$ Phasenraum der Materie, $v$ is Pneuma, $u$ is Soma [61, Sec. 35].
Let ${\mathcal{X}}^{J}_{1}\approx{\mathcal{X}}_{1}\times\mathbb{R}^{2}$ be the Siegel–Jacobi upper half-plane, where ${\mathcal{X}}_{1}=\{v\in\mathbb{C}|~{}y:=\operatorname{Im}v>0\}$ is the Siegel upper half-plane, and $\tilde{{\mathcal{X}}}^{J}_{1}\approx{\mathcal{X}}^{J}_{1}\times\mathbb{R}$ denotes the extended Siegel–Jacobi upper half-plane. Then:
Lemma 1.
a) The action $G^{J}(\mathbb{R})_{0}\times{\mathcal{X}}^{J}_{1}\rightarrow{\mathcal{X}}^{J}_{1}$ is given by
(2.4) $(M,X)\times(v^{\prime},u^{\prime})=(v_{1},u_{1}),\emph{\text{~{}where~{}}}v_{1}=\frac{av^{\prime}+b}{cv^{\prime}+d},~{}u_{1}=\frac{u^{\prime}+\lambda u^{\prime}+\mu}{cu^{\prime}+d}.$
b) If $u^{\prime}=p^{\prime}v^{\prime}+q^{\prime}$, $v^{\prime}=x^{\prime}+\operatorname{i}y^{\prime}$ as in (2.3), then the action
(2.5) $(M,X)\times(x^{\prime},y^{\prime},p^{\prime},q^{\prime})=(x_{1},y_{1},p_{1},q_{1})$
is given by the formula
(2.6) $x_{1}+\operatorname{i}y_{1}=\frac{(ax^{\prime}+b)(cx^{\prime}+d)+acy^{\prime 2}+\operatorname{i}y^{\prime}}{(cx^{\prime}+d)^{2}+(cy^{\prime})^{2}},$
and
(2.7) $(p_{1},q_{1})=(p,q)+(p^{\prime},q^{\prime})\left(\begin{array}[]{cc}a&b\\ c&d\end{array}\right)^{-1}=(p+dp^{\prime}-cq^{\prime},q-bp^{\prime}+aq^{\prime}).$
c) The action $G^{J}_{1}(\mathbb{R})\times\tilde{{\mathcal{X}}}^{J}_{1}\rightarrow\tilde{{\mathcal{X}}}^{J}_{1}$ is given by
(2.8) $\begin{split}&(M,X,\kappa)\times(v^{\prime},z^{\prime},\kappa^{\prime})=(v_{1},z_{1},\kappa_{1}),\\ &(M,X,\kappa)\times(x^{\prime},y^{\prime},p^{\prime},q^{\prime},\kappa^{\prime})=(x_{1},y_{1},p_{1},q_{1},\kappa_{1}),\\ &\kappa_{1}=\kappa+\kappa^{\prime}+\lambda q^{\prime}-\mu p^{\prime},~{}(p^{\prime},q^{\prime})=(\frac{\rho^{\prime}}{y^{\prime}},\xi^{\prime}-\frac{x^{\prime}}{y^{\prime}}\rho^{\prime}),~{}(\lambda,\mu)=(p,q)M\end{split}$
and (2.6), (2.7).
Proposition 1 is an improved version of [22, (4.38), (5.8)], [25, (28),  (29)], [27, Proposition 2.1], [29, Proposition 2], [30, Proposition 2], [17, (18),(19)].
Below $(w,z)\in({\mathcal{D}}_{1},\mathbb{C})$, $(v,u)\in({\mathcal{X}}_{1},\mathbb{C})$, and the parameters $k$ and $\nu$ come from representation theory of the Jacobi group: $k$ indexes the positive discrete series of ${\rm SU}(1,1)$, $2k\in\mathbb{N}$, while $\nu>0$ indexes the representations of the Heisenberg group [15]. $\rm{FC}$ is an abbreviation for the fundamental conjecture for homogeneous Kähler  manifolds [86], see also [52], [55].
Proposition 1.
Perelomov’s coherent state vectors associated to the group $G^{J}_{1}$ are defined as
(2.9) $e_{z,w}:=e^{\sqrt{\mu}z{{\mbox{\boldmath{$a$}}}}^{\dagger}+w{{\mbox{\boldmath{$K$}}}}_{+}}e_{0},~{}z\in\mathbb{C},~{}|w|<1,$
and the reproducing kernel $K=K(\bar{z},\bar{w},z,w)$ is
(2.10) $K=\!(e_{z,w},e_{z,w})\!=\!(1\!-\!w\bar{w})^{-2k}\exp{\nu\frac{2z\bar{z}\!+\!z^{2}\bar{w}\!+\!\bar{z}^{2}w}{2(1-w\bar{w})}},z,w\in\mathbb{C},|w|<1.$
a) The Kähler two-form on ${\mathcal{D}}^{J}_{1}$, invariant to the action of $G^{J}_{1}=\rm{SU}(1,1)\ltimes\mathbb{C}$, is
(2.11) $-\operatorname{i}\omega_{{\mathcal{D}}^{J}_{1}}(w,z)\!=\!\frac{2k}{P^{2}}\operatorname{d}w\wedge\operatorname{d}\bar{w}\!+\!\nu\frac{{\mathcal{A}}\wedge\bar{{\mathcal{A}}}}{P},P:=1-|w|^{2},{\mathcal{A}}={\mathcal{A}}(w,z):=\operatorname{d}z\!+\!\bar{\eta}\operatorname{d}w.$
We have the change of variables $FC:(w,z)\rightarrow(w,\eta,\bar{\eta})$
(2.12) ${\rm FC}\colon\ z=\eta-w\bar{\eta},\qquad{\rm FC}^{-1}\colon\ \eta=\frac{z+\bar{z}w}{P},$
and
(2.13) ${\rm FC}\colon\ {\mathcal{A}}(w,z)\rightarrow{\mathcal{A}}(w,\eta,\bar{\eta}):=\operatorname{d}\eta-w\operatorname{d}\bar{\eta},$
(2.14a) $-\operatorname{i}\omega_{{\mathcal{D}}^{J}_{1}}(w,\eta)$ $=-\operatorname{i}{\rm{FC}}^{*}(\omega_{{\mathcal{D}}^{J}_{1}}(w,z))=\frac{2k}{P^{2}}\operatorname{d}w\wedge\operatorname{d}\bar{w}+\nu\operatorname{d}\eta\wedge\operatorname{d}\bar{\eta},$
(2.14b) $\omega_{{\mathcal{D}}^{J}_{1}}(\alpha,\beta,q,p)$ $=4k\frac{\operatorname{d}\alpha\wedge\operatorname{d}\beta}{(1-\alpha^{2}-\beta^{2})^{2}}+2\nu\operatorname{d}q\wedge\operatorname{d}p,$
where
(2.15) $w=\alpha+\operatorname{i}\beta,\alpha,~{}\beta\in\mathbb{R},\quad\eta=q+\operatorname{i}p,~{}p,q\in\mathbb{R}.$
Also with (2.10) and (2.12) we have
(2.16) $B(w,\eta-w\bar{\eta})=(1-w\bar{w})^{-2k}\exp\nu\left[\eta\bar{\eta}-\frac{\bar{w}\eta^{2}+w\bar{\eta}^{2}}{2}\right].$
With a formula similar to [15, (7.18)] applied to (2.16), we get
(2.17a) $-\operatorname{i}\omega(w,\eta)$ $=\frac{2k}{(1-w\bar{w})^{2}}\operatorname{d}w\wedge\operatorname{d}\bar{w}+\nu[\operatorname{d}\eta\wedge\operatorname{d}\bar{\eta}-\bar{\eta}\operatorname{d}w\wedge\operatorname{d}\bar{\eta}+\eta\operatorname{d}\bar{w}\wedge\operatorname{d}\eta],$
(2.17b) $\omega(\alpha,\beta,q,p)$ $=4k\frac{\operatorname{d}\alpha\wedge\operatorname{d}\beta}{(1-\alpha^{2}-\beta^{2})^{2}}+2\nu\operatorname{d}q\wedge\operatorname{d}p$
$+2\nu[\operatorname{d}q\wedge(p\operatorname{d}\alpha-q\operatorname{d}\beta)+\operatorname{d}p\wedge(p\operatorname{d}\beta+q\operatorname{d}\alpha)],$
and equation (2.17b) is different of (2.14b).
In (2.16) we make the change of coordinates $w\rightarrow v$ (2.28b) and $\eta=q+\operatorname{i}p$, we get
$\bar{\eta}^{2}w=\frac{(q^{2}-p^{2}-2\operatorname{i}qp)(x^{2}+y^{2}-1-2\operatorname{i}x)}{N},$
and finally we get
(2.18) $f(x,y,q,p)=-2k\log\frac{4y}{N}+\nu F,\quad F=\frac{2}{N}[(y+1)q^{2}+(x^{2}+y^{2}+y)p^{2}+2qpx].$
If in (2.17a) we make the change of variables $w\rightarrow v$ (2.24a), we get the Kähler  two-form
(2.19) $-\operatorname{i}\omega(v,\eta)=\frac{k}{2y^{2}}\operatorname{d}v\wedge\operatorname{d}\bar{v}+\nu\{-2\operatorname{i}[\frac{\bar{\eta}}{(v+\operatorname{i})^{2}}\operatorname{d}v\wedge\operatorname{d}\bar{\eta}+\frac{\eta}{(\bar{v}-\operatorname{i})^{2}}\operatorname{d}\bar{v}\wedge\operatorname{d}\eta]+\operatorname{d}\eta\wedge\operatorname{d}\bar{\eta}\},$
or the symplectic two-form
$\omega_{{\mathcal{X}}^{J}_{1}}(x,y,q,p)$ $=\frac{k}{y^{2}}\operatorname{d}x\wedge\operatorname{d}y\!+\!\frac{4\nu}{N^{2}}\{[q(x^{2}\!-\!(y+1)^{2})\!-\!2px(y+1)](\operatorname{d}x\wedge\operatorname{d}q+\operatorname{d}y\wedge\operatorname{d}p)$
$+[2qx(y+1)+p(x^{2}-(y+1)^{2})](-\operatorname{d}x\wedge\operatorname{d}p+\operatorname{d}y\wedge\operatorname{d}q)\}+2\nu\operatorname{d}q\wedge\operatorname{d}p.$
Note that (1) is different from (2.25b) and (2.49). The matrix of the balanced metric $h=h(\varsigma)$, $\varsigma:=(z,w)\in\mathbb{C}\times{\mathcal{D}}_{1}$ associated to the Kähler two-form (2.11) is
(2.21) $h(\varsigma)=\left(\begin{array}[]{cc}h_{z\bar{z}}&h_{z\bar{w}}\\ h_{w\bar{z}}&h_{w\bar{w}}\end{array}\right)=\left(\begin{array}[]{cc}\frac{\mu}{P}&\mu\frac{\eta}{P}\\ \mu\frac{\bar{\eta}}{P}&\frac{2k}{P^{2}}+\mu\frac{|\eta|^{2}}{P}\end{array}\right).$
The inverse of the matrix (2.21) reads
(2.22) $h^{-1}(\varsigma)=\left(\begin{array}[]{cc}h^{z\bar{z}}&h^{z\bar{w}}\\ h^{w\bar{z}}&h^{w\bar{w}}\end{array}\right)=\left(\begin{array}[]{cc}\frac{P}{\mu}+\frac{P^{2}|\eta|^{2}}{2k}&-\frac{P^{2}\eta}{2k}\\ -\frac{P^{2}\bar{\eta}}{2k}&\frac{P^{2}}{2k}\end{array}\right).$
b) The second partial Cayley transform $\Phi_{1}:{\mathcal{D}}^{J}_{1}\rightarrow{\mathcal{X}}^{J}_{1}$ and
(2.23) $\Phi_{1}:={\rm{FC}}_{1}\circ\Phi:(w,z)\rightarrow(v=x+\operatorname{i}y,\eta=q+\operatorname{i}p)$
and its inverse $\Phi_{1}^{-1}:(v,\eta)\rightarrow(w,z)$ are given by
(2.24a) $\Phi_{1}:$ $~{}w=\frac{v-\operatorname{i}}{v+\operatorname{i}},\quad z=\eta-\bar{\eta}\frac{v-i}{v+\operatorname{i}}=2\operatorname{i}\frac{pv+q}{v+\operatorname{i}},$
(2.24b) $\Phi_{1}^{-1}:$ $~{}v=\operatorname{i}\frac{1+w}{1-w},\quad\eta=\frac{(1+\operatorname{i}\bar{v})(z-\bar{z})+v(\bar{v}-\operatorname{i})(z+\bar{z})}{2\operatorname{i}(\bar{v}-v)}=\frac{z+\bar{z}w}{P}.$
Introducing the second partial Cayley transform (2.24a) into the Kähler two-form (2.11) on ${\mathcal{D}}^{J}_{1}$, we get the symplectic two-form (1) on the Siegel-Jacobi upper half-plane $(v,\eta)$, $\operatorname{Im}v>0$
(2.25a) $-\operatorname{i}\omega_{{\mathcal{X}}^{J}_{1}}(v,\bar{v},\eta,\bar{\eta})$ $=\frac{8k}{P^{2}}\frac{\operatorname{d}v\wedge\operatorname{d}\bar{v}}{N^{2}}+\nu\operatorname{d}\eta\wedge\operatorname{d}\bar{\eta}=-\frac{2k}{(\bar{v}-v)^{2}}\operatorname{d}v\wedge\operatorname{d}\bar{v}+\eta\wedge\operatorname{d}\bar{\eta},$
(2.25b) $\omega_{{\mathcal{X}}^{J}_{1}}(x,y,q,p)$ $=\frac{k}{y^{2}}\operatorname{d}x\wedge\operatorname{d}y+2\nu\operatorname{d}q\wedge\operatorname{d}p,$
(2.26) $N:=|v+\operatorname{i}|^{2}=x^{2}+(y+1)^{2},$
(2.27) $P=4\frac{y}{N}.$
c) Using the partial Cayley transform $\Phi^{-1}:{\mathcal{D}}^{J}_{1}\rightarrow{\mathcal{X}}^{J}_{1},~{}(w,z)\rightarrow(v,u)$ and its inverse
(2.28a) $\Phi^{-1}:v$ $=\operatorname{i}\frac{1+w}{1-w},~{}~{}u=\frac{z}{1-w},~{}~{}w,z\in\mathbb{C},~{}|w|<1,$
(2.28b) $\Phi:w$ $=\frac{v-\operatorname{i}}{v+\operatorname{i}},~{}~{}z=2\operatorname{i}\frac{u}{v+\operatorname{i}},~{}~{}v,u\in\mathbb{C},~{}\operatorname{Im}v>0,$
we obtain
(2.29) ${\mathcal{A}}\left(\frac{v-\operatorname{i}}{v+\operatorname{i}},\frac{2\operatorname{i}u}{v+\operatorname{i}}\right)=\frac{2\operatorname{i}}{v+\operatorname{i}}{\mathcal{B}}(v,u),$
where
(2.30) ${\mathcal{B}}(v,u):={\rm d}u-r{\rm d}v,~{}r:=\frac{u-\bar{u}}{v-\bar{v}}.$
The Berndt–Kähler’s two-form (symplectic two-form) invariant to the action of $G^{J}(\mathbb{R})_{0}$ =SL(2,)absentleft-normal-factor-semidirect-productnormal-$=\rm{SL}(2,\mathbb{R})\ltimes\mathbb{C}$, is (2.31a) ((2.31c))
(2.31a) $-\operatorname{i}\omega_{{\mathcal{X}}^{J}_{1}}(v,u)$ $=-\frac{2k}{(\bar{v}-v)^{2}}\operatorname{d}v\wedge\operatorname{d}\bar{v}+\frac{2\nu}{\operatorname{i}(\bar{v}-v)}{\mathcal{B}}\wedge\bar{{\mathcal{B}}}$
(2.31b) $=\frac{1}{y}\{(\frac{k}{2y}+\nu r^{2})\operatorname{d}v\wedge\operatorname{d}\bar{v}+\nu[\operatorname{d}u\wedge\operatorname{d}\bar{u}-r(\operatorname{d}v\wedge\operatorname{d}\bar{u}-cc)]\},$
(2.31c) $\omega(x,y,m,n)$ $=\frac{k}{y^{2}}\operatorname{d}x\wedge\operatorname{d}y+\frac{2\nu}{y}(\operatorname{d}m-r\operatorname{d}x)\wedge(\operatorname{d}n-r\operatorname{d}y)$
(2.31d) $=(\frac{k}{y^{2}}\!+\!2\nu\frac{r^{2}}{y})\operatorname{d}x\wedge\operatorname{d}y\!+\!2\frac{\nu}{y}[\operatorname{d}m\wedge\operatorname{d}n\!+\!r(\operatorname{d}y\wedge\operatorname{d}m\!-\!\operatorname{d}x\wedge\operatorname{d}n)],$
(2.32) $u=m+\operatorname{i}n,\quad m,n\in\mathbb{R},\quad r=\frac{n}{y}$
With (2.10) and (2.28b) we get
(2.33) $K(\frac{v-\operatorname{i}}{v+\operatorname{i}},\frac{2\operatorname{i}u}{v+\operatorname{i}})=\left[\frac{|v+\operatorname{i}|^{2}}{2\operatorname{i}(\bar{v}-v)}\right]^{2k}\exp\frac{2\nu}{|v+\operatorname{i}|^{2}}[|u|^{2}-\frac{(u\bar{v}-\bar{u}v)^{2}+(\bar{u}-u)^{2}}{2\operatorname{i}(\bar{v}-v)}].$
With the change of variables ${\rm FC}_{1}\colon(v,u)\rightarrow(v,\eta)$
[…]
in (2.31a) we get
$r=p,\quad m=px+q,\quad n=py,$
$B(v,\bar{v},\eta,\bar{\eta})=\frac{1}{2\operatorname{i}}[(v+\operatorname{i})\operatorname{d}\eta-(v-\operatorname{i})\operatorname{d}\bar{\eta}],$
and finally we regain (1).
The matrix corresponding to the balanced metric (1) associated with the Kähler  two -form (2.31a) reads
(2.35) $h(v,u)=\left(\begin{array}[]{cc}h_{v\bar{v}}&h_{v\bar{u}}\\ \bar{h}_{v\bar{u}}&h_{u\bar{u}}\end{array}\right)=\left(\begin{array}[]{cc}\frac{k}{2y^{2}}+\nu\frac{r^{2}}{y}&-\nu\frac{r}{y}\\ -\nu\frac{r}{y}&\frac{\nu}{y}\end{array}\right),~{}y:=\frac{v-\bar{v}}{2\operatorname{i}},$
and we also have
(2.36) $h^{-1}(v,u)=\left(\begin{array}[]{cc}h^{v\bar{v}}&h^{v\bar{u}}\\ \bar{h}^{u\bar{v}}&h^{u\bar{u}}\end{array}\right)=\left(\begin{array}[]{cc}\frac{2y^{2}}{k}&\frac{2y^{2}r}{k}\\ \frac{2y^{2}r}{k}&\frac{y}{\nu}+2\frac{r^{2}y^{2}}{k}\end{array}\right).$
d) If we apply the change of coordinates ${\mathcal{D}}^{J}_{1}\ni(v,u)\rightarrow(x,y,p,q)\in{\mathcal{X}}^{J}_{1}$ (2.3),
then
(2.37) ${\mathcal{B}}(v,u)=\operatorname{d}u-p\operatorname{d}v,$
(2.38) ${\mathcal{B}}(v,u)={\mathcal{B}}(x,y,p,q):=F\operatorname{d}t={\mathcal{F}}=v\operatorname{d}p+\operatorname{d}q=(x+\operatorname{i}y)\operatorname{d}p+\operatorname{d}q,F:=\dot{p}v+\dot{q},$
and we regain (2.25b).
e) The two-parameter balanced metric on the Siegel–Jacobi upper half-plane ${\mathcal{X}}^{J}_{1}$ associated to the Kähler  two-form (2.31a) is
(2.39a) $\!\!\operatorname{d}s^{2}_{{\mathcal{X}}^{J}_{1}}(x,y,p,q)\!$ $=\!\alpha\frac{\operatorname{d}x^{2}\!+\!\operatorname{d}y^{2}}{y^{2}}\!+\!\frac{\gamma}{y}(S\operatorname{d}p^{2}\!+\!\operatorname{d}q^{2}+2x\operatorname{d}p\operatorname{d}q)$
(2.39b) $=\!\alpha\frac{\operatorname{d}x^{2}\!+\!\operatorname{d}y^{2}}{y^{2}}\!+\!\frac{\gamma}{y}(A^{2}\!+\!\!B^{2}),$
where
(2.40) $\alpha:=k/2,\quad\gamma:=\nu,\quad S:=x^{2}+y^{2},A\!:=\!\operatorname{Re}{\mathcal{F}}\!=\!x\operatorname{d}p\!+\!\operatorname{d}q,B\!:=\!\operatorname{Im}{\mathcal{F}}\!=\!p\operatorname{d}y.$
The metric matrix associated with (1) is
$g_{{{\mathcal{X}}}^{J}_{1}}\!=\!\left(\begin{array}[]{cccc}g_{xx}&0&0&0\\ 0&g_{yy}&0&0\\ 0&0&g_{pp}&g_{pq}\\ 0&0&g_{qp}&g_{qq}\end{array}\right),\!\begin{array}[]{cc}\qquad g_{xx}\!=\frac{\alpha}{y^{2}},&\!g_{yy}\!=\!\frac{\alpha}{y^{2}};\\ g_{pq}\!=\!\gamma\frac{x}{y},&g_{pp}\!=\!\gamma\frac{S}{y},\quad g_{qq}\!=\!\frac{\gamma}{y}.\end{array}$
Below we reproduce the Comment 5.5 in the first reference [27] with some completions:
Comment 1.
Berndt [39, p 8] considered the closed two-form $\Omega=\operatorname{d}\bar{\operatorname{d}}f^{\prime}$ of Siegel–Jacobi upper half-plane ${\mathcal{X}}^{J}_{1}$, $G^{J}(\mathbb{R})_{0}$-invariant to the action (2.4), obtained from the Kähler potential
(2.41) $f^{\prime}(\tau,z)=c_{1}\log(\tau-\bar{\tau})-\operatorname{i}c_{2}\frac{(z-\bar{z})^{2}}{\tau-\bar{\tau}},~{}c_{1},c_{2}>0.$
Formula (2.41) is presented by Berndt as “communicated to the author by Kähler”. Also in [39, p 8] is given our equation (5.21a) in first reference [27], while our present equation (1) corrects two printing errors in Berndt’s paper.
Later, in [61, §  36], reproduced also in [62], Kähler  argues how to choose the potential as in (2.41), see also [61, (9) §  37], where $c_{1}=-\frac{k}{2}$, $c_{2}=\operatorname{i}\nu\pi$, i. e.
(2.42) $f^{\prime}(\tau,z)=-\frac{k}{2}\log\frac{\tau-\bar{\tau}}{2\operatorname{i}}-\operatorname{i}\pi\nu\frac{(z-\bar{z})^{2}}{\tau-\bar{\tau}}.$
Once the Kähler  potential (2.42) is known, we apply the recipe (3.2b)
$-\operatorname{i}\omega_{{\mathcal{X}}^{J}_{1}}(\tau,z)=f^{\prime}_{\tau\bar{\tau}}\operatorname{d}\tau\wedge\operatorname{d}\bar{\tau}+f^{\prime}_{\tau\bar{z}}\operatorname{d}\tau\wedge\operatorname{d}\bar{z}-\bar{f}^{\prime}_{\tau\bar{z}}\operatorname{d}\bar{\tau}\wedge\operatorname{d}z+f^{\prime}_{z\bar{z}}\operatorname{d}z\wedge\operatorname{d}\bar{z}.$
The metric (8) in [61] differs from the metric (1) by a factor of two, since the Hermitian metric used by Kähler  is $\operatorname{d}s^{2}=2g_{i\bar{j}}\operatorname{d}z_{i}\otimes\operatorname{d}\bar{z}_{j}$. If in (2.42) we take $k/2\rightarrow k$, we have
$f^{\prime}_{\tau}$ $=-k\frac{1}{\tau-\bar{\tau}}+\operatorname{i}\pi\nu\frac{(z-\bar{z})^{2}}{(\tau-\bar{\tau})^{2}},~{}f^{\prime}_{\tau\bar{\tau}}=-k\frac{1}{(\tau-\bar{\tau})^{2}}+2\operatorname{i}\pi\nu\frac{(z-\bar{z})^{2}}{(\tau-\bar{\tau})^{3}},$
$f^{\prime}_{\tau\bar{z}}$ $=-2\operatorname{i}\pi\nu\frac{z-\bar{z}}{(\tau-\bar{\tau})^{2}},~{}f^{\prime}_{z}=-2\operatorname{i}\pi\nu\frac{z-\bar{z}}{\tau-\bar{\tau}},~{}f^{\prime}_{z\bar{z}}=2\operatorname{i}\pi\nu\frac{1}{\tau-\bar{\tau}},$
and we get (2.31a). Relation (2.31a) has been obtained by Berndt [38, p 30], where the denominator of the first term is misprinted as $v-\bar{v}$ (or $\tau-\bar{\tau}$ in our notations). Equation (1) appears also in [38, p 30] and [40, p 62].
We denote in (2.42) $(\tau,z)$ with $(v,u)$ as in [15, (9.16)]. Indeed, we make successively the transformations: the partial Cayley transform, $\Phi:~{}(w,z)\rightarrow(v,u)$ (1), a holomorphic transform, and we get (2.44a), then we apply the $FC_{1}$ transform $(v,u)\rightarrow(v,\eta)$ (2.34), a non-holomorphic transform, to obtain (2.44b), and finally we make the symplectic transform $(w,z)\rightarrow(x,y,q,p)$ with the result (2.44c)
(2.44a) $\log K(v,u)$ $=-\frac{k}{2}\log\frac{v-\bar{v}}{2\operatorname{i}}-\operatorname{i}\nu\frac{(u-\bar{u})^{2}}{v-\bar{v}},$
(2.44b) $\log K(v,\eta)$ $=-\frac{k}{2}\log\frac{v-\bar{v}}{2\operatorname{i}}+\frac{\operatorname{i}\nu}{4}(\eta-\bar{\eta})^{2}(v-\bar{v}),$
(2.44c) $\log K(x,y,q,p)$ $=-\frac{k}{2}\log y+2\nu yp^{2}.$
Note that (2.44a) is different of (2.33).
The metric associated to the Kähler two-form (2.25b) is
(2.45) $\operatorname{d}s^{2}(x,y,q,p)=\frac{k}{2y^{2}}(\operatorname{d}x^{2}+\operatorname{d}y^{2})+\nu(\operatorname{d}q^{2}+\operatorname{d}p^{2}).$
The metric corresponding to the Kähler two-form (2.44a) is
(2.46a) $\operatorname{d}s^{2}(x,y,n,m)$ $\!=\!(\frac{k}{2}+\nu\frac{n^{2}}{y})\frac{\operatorname{d}x^{2}\!+\!\operatorname{d}y^{2}}{y^{2}}\!+\!\frac{\nu}{y}[\operatorname{d}n^{2}\!+\!\operatorname{d}m^{2}-2r(\operatorname{d}m\operatorname{d}x\!+\!\operatorname{d}n\operatorname{d}y)]$
(2.46b) $=\frac{k}{2}\frac{\operatorname{d}x^{2}+\operatorname{d}y^{2}}{y^{2}}+\frac{\nu}{y}[(r\operatorname{d}x-\operatorname{d}m)^{2}+(r\operatorname{d}y-\operatorname{d}n)^{2}],~{}r=\frac{n}{y}.$
The metric (1) corresponds to the Kähler potential (2.44a)
(2.47) $f"(v,u)=-2k\log\frac{v-\bar{v}}{2\operatorname{i}}-\operatorname{i}\nu\frac{(u-\bar{u})^{2}}{v-\bar{v}}$
instead of (2.42).
Equation (2.44c) was presented in [15, (9.20)].
In [36, (4.3)], see also [36, Proposition 4.1], we have presented a generalization of (2.44c) for ${\mathcal{X}}^{J}_{n}$, obtained by Takase in [85, §9].
Yang calculated in [92] the metric on ${\mathcal{X}}^{J}_{n}$, invariant to the action of $G^{J}_{n}(\mathbb{R})_{0}$. The equivalence of the metric of Yang with the metric obtained via CS on ${\mathcal{D}}^{J}_{n}$ and then transported to ${\mathcal{X}}^{J}_{n}$ via the partial Cayley transform $(v,u)\rightarrow(v,\eta)$ (1) is underlined in [19]. In particular, the metric (5.21b) in the first reference [27] appears in [92, p 99] for the particular values $c_{1}=1$, $c_{2}=4$. See also [91, 93, 94].
Remark 1.
In formula (2.44a) we make the change of variables $FC_{1}$ (2.34) and we get (2.44b). We apply to (2.44b) [15, (7.18)] to calculate
$-\operatorname{i}\omega(v,\eta)=h_{v\bar{v}}\operatorname{d}v\wedge\operatorname{d}\bar{v}+h_{v\bar{\eta}}\operatorname{d}v\wedge\operatorname{d}\bar{\eta}-\bar{h}_{v\bar{\eta}}\operatorname{d}\bar{v}\wedge\operatorname{d}\eta+h_{\eta\bar{\eta}}\operatorname{d}\eta\wedge\operatorname{d}\bar{\eta}.$
The associated matrix
(2.48) $h=\left(\begin{array}[]{cc}h_{v\bar{v}}&h_{v\bar{\eta}}\\ h_{\eta\bar{v}}&h_{\eta\bar{\eta}}\end{array}\right)=\left(\begin{array}[]{cc}\frac{k}{8y^{2}}&\nu p\\ \nu p&\nu y\end{array}\right)$
is hermitian and we have
(2.49) $\omega_{{\mathcal{X}}^{J}_{1}}(x,y,q,p)=\frac{k}{4}\frac{\operatorname{d}x\wedge\operatorname{d}y}{y^{2}}+2\nu[p(\operatorname{d}x\wedge\operatorname{d}p+\operatorname{d}q\wedge\operatorname{d}y)+y\operatorname{d}q\wedge\operatorname{d}p].$
(2.49) is different of (2.25b) obtained introducing (2.24a) into (2.11).
Proof.
We get from (2.44b)
(2.50a) $h_{v}$ $=-\frac{k}{2}\frac{1}{v-\bar{v}}+\frac{\operatorname{i}\nu}{4}(\eta-\bar{\eta})^{2},$
(2.50b) $h_{v\bar{v}}$ $=-\frac{k}{2}(v-\bar{v})^{-2}=\frac{k}{8}\frac{1}{y^{2}},$
(2.50c) $h_{v\bar{\eta}}$ $=-\operatorname{i}\frac{\nu}{2}(\eta-\bar{\eta})=\nu p,$
(2.50d) $h_{\eta}$ $=\frac{\operatorname{i}\nu}{2}(\eta-\bar{\eta})(v-\bar{v}),$
(2.50e) $h_{\eta\bar{\eta}}$ $=\nu y,$
and we get (2.48) which is Hermitian. The conditions (3.4) that the metric associated to (2.49) be Kähler  are met. ∎
We have also obtained invariant metric to the action of the Jacobi group $G^{J}_{1}(\mathbb{R})$ on the extended Siegel-Jacobi upper half-plane $\tilde{{\mathcal{X}}}^{J}_{1}$ [27, Proposition 5.6, (5.25), (5.26)], see also [4, Proposition 4, (69) ]
Proposition 2.
The three-parameter metric of the extended Siegel-Jacobi upper half-plane $\tilde{{\mathcal{X}}}^{J}_{1}$ expressed in the S-coordinates $(x,y,p,q,\kappa)$, left-invariant with respect to the action of the Jacobi group $G^{J}_{1}(\mathbb{R})$, is given as
(2.51) $\begin{split}{\rm d}s^{2}_{\tilde{{\mathcal{X}}}^{J}_{1}}(x,y,p,q,\kappa)&\!=\!{\rm d}s^{2}_{{\mathcal{X}}^{J}_{1}}(x,y,p,q)+\lambda^{2}_{6}(p,q,\kappa)\\ &\!=\!\frac{\alpha}{y^{2}}\big{(}{\rm d}x^{2}+{\rm d}y^{2}\big{)}+\left(\frac{\gamma}{y}S+\delta q^{2}\right){\rm d}p^{2}+\left(\frac{\gamma}{y}+\delta p^{2}\right){\rm d}q^{2}+\delta{\rm d}\kappa^{2}\\ &\!+2\left(\gamma\frac{x}{y}-\delta pq\right){\rm d}p{\rm d}q+2\delta(q{\rm d}p{\rm d}\kappa-p{\rm d}q{\rm d}\kappa),\end{split}$
where [27, (5.15f), (5.17)]
$\lambda_{6}=\sqrt{\delta}(\operatorname{d}\kappa-p\operatorname{d}q+q\operatorname{d}p),$
and $S$ was defined in (2.40).
The metric matrix associated to the metric (2.51) is
(2.52) $g_{\tilde{{\mathcal{X}}}^{J}_{1}}\!=\!\left(\begin{array}[]{ccccc}g_{xx}&0&0&0&0\\ 0&g_{yy}&0&0&0\\ 0&0&g_{pp}&g_{pq}&g_{p\kappa}\\ 0&0&g_{qp}&g_{qq}&g_{q\kappa}\\ 0&0&g_{\kappa p}&g_{\kappa q}&g_{\kappa\kappa}\end{array}\right),\!\begin{array}[]{cc}g_{xx}\!=\frac{\alpha}{y^{2}},&\!g_{yy}\!=\!\frac{\alpha}{y^{2}},\\ g_{pq}\!=\!\gamma\frac{x}{y}-\delta pq,&~{}g_{p\kappa}\!=\!\delta q,g_{q\kappa}\!=\!-\delta p,\\ g_{pp}\!=\!\gamma\frac{S}{y}+\delta q^{2},&g_{qq}\!=\!\frac{\gamma}{y}+\delta p^{2},g_{\kappa\kappa}\!=\!\delta.\end{array}$
The extended Siegel–Jacobi upper half-plane $\tilde{{\mathcal{X}}}^{J}_{1}$ does not admit an almost contact structure $(\Phi,\xi,\eta)$ with a contact form $\eta=\lambda_{6}$ and Reeb vector $\xi=\operatorname{Ker}(\eta)$.
Now some results of Proposition 1 are extended from ${\mathcal{D}}^{J}_{1}$ and ${\mathcal{X}}^{J}_{1}$ to ${\mathcal{D}}^{J}_{n}$, respectively ${\mathcal{X}}^{J}_{n}$. Below $k$, $2k\in\mathbb{N}$ indexes the holomorphic discrete series of ${\mbox{\rm{Sp}}(n,\mathbb{R})}$ and $\nu>0$ indexes the representations of the Heisenberg group. Parts of the following Proposition are taken from [30, Proposition 3, Theorem 1], see also [19, Proposition 3], [26, Theorem 3.2]:
Proposition 3.
a) The Kähler two-form on $n(n+3)$-dimensional ${\mathcal{D}}^{J}_{n}$, invariant to the action of $(G^{J}_{n})_{0}$, is
(2.53a) $\!-\!\operatorname{i}\!\omega_{{\mathcal{D}}^{J}_{n}}(W,z)$ $\!=\!\frac{k}{2}\operatorname{tr}(B\wedge\bar{B})\!+\!\nu\operatorname{tr}(A^{t}\bar{M}\wedge\bar{A}),\!A(W,z)\!:=\!\operatorname{d}z^{t}\!+\!\operatorname{d}W\bar{\eta},W\in{\mathcal{D}}_{n},$
(2.53b) $B(W)$ $:=M\operatorname{d}W,~{}M\!:=\!(\mathbb{1}_{n}-W\bar{W})^{-1},~{}z\in M(1,n,\mathbb{C}),~{}\eta\in M(n,1,\mathbb{C}),$
b) Using the partial Cayley transform
(2.54a) $\Phi^{-1}$ $:v=\operatorname{i}(\mathbb{1}_{n}-W)^{-1}(\mathbb{1}_{n}+W);~{}u^{t}=(\mathbb{1}_{n}-W)^{-1}z^{t},~{}W\in{\mathcal{D}}_{n},~{}v\in{\mathcal{X}}_{n};$
(2.54b) $\Phi$ $:W=(v-\operatorname{i}\mathbb{1}_{n})^{-1}(v+\operatorname{i}\mathbb{1}_{n}),~{}z^{t}=2\operatorname{i}(v+\operatorname{i}\mathbb{1}_{n})^{-1}u^{t},~{}z,u\in M(1,n,\mathbb{C}),$
we get from the Kähler  two-form on ${\mathcal{X}}^{J}_{n}$ depending on two parameters, invariant to the action of $G^{J}_{n}(\mathbb{R})_{0}$:
(2.55) $-\operatorname{i}\omega_{{\mathcal{X}}^{J}_{n}}(v,u)=\frac{k}{2}\operatorname{tr}(H\wedge\bar{H})+\frac{2\nu}{\operatorname{i}}\operatorname{tr}(G^{t}D\wedge\bar{G}),\quad D:=(\bar{v}-v)^{-1},~{}H:=D\operatorname{d}v.$
where
(2.56) $G^{t}(v,u)=\operatorname{d}u-p\operatorname{d}v,$
and
(2.57) $G^{t}(v,u)=G^{t}(x,y,p,q)=\operatorname{d}pv+\operatorname{d}q=\operatorname{d}p(x+\operatorname{i}y)+\operatorname{d}q.$
c) Let $M(n,\mathbb{C})\ni v=x+\operatorname{i}y$ be a symmetric positive definite matrix and $p,q\in M(n,1,\mathbb{C})$. The three parameter metric on $\tilde{{\mathcal{X}}}^{J}_{n}$, invariant to the $G^{J}_{n}(\mathbb{R})$ action is
(2.58) $\begin{split}\operatorname{d}s_{\tilde{{\mathcal{X}}}^{J}_{n}}^{2}(x,y,p,q,\kappa)&=\operatorname{d}s_{{\mathcal{X}}^{J}_{n}}^{2}(x,y,p,q)+\lambda_{6}^{2}\\ &=\alpha\operatorname{tr}[(y^{-1}\operatorname{d}x)^{2}+(y^{-1}\operatorname{d}y)^{2}]\\ &+\gamma[\operatorname{d}p(xy^{-1}x+yy^{-1}y)\operatorname{d}p^{t}+\operatorname{d}qy^{-1}\operatorname{d}q^{t}+2\operatorname{d}pxy^{-1}\operatorname{d}q^{t}]\\ &+\delta(\operatorname{d}\kappa-p\operatorname{d}q^{t}+q\operatorname{d}p^{t})^{2}.\end{split}$
3. Berry phase on Kähler  manifolds
3.1. Balanced metric
The starting point in Perelomov’s approach to coherent states (CS) is the triplet $(G,\pi,{\mathfrak{H}})$, where $\pi$ is a unitary, irreducible representation of the Lie group $G$ on a separable complex Hilbert space ${\mathfrak{H}}$ [79].
Two types of CS-vectors belonging to ${\mathfrak{H}}$ are locally defined on $M=G/H$: the normalized (un-normalized) CS-vector $\underline{e}_{x}$ (respectively, $e_{z}$) [7, §6, Remark 4, (6.25)]
(3.1) $\underline{e}_{x}=\exp(\sum_{\phi\in\Delta^{+}}x_{\phi}{{\mbox{\boldmath{$X$}}}}^{+}_{\phi}-{\bar{x}}_{\phi}{{\mbox{\boldmath{$X$}}}}^{-}_{\phi})e_{0},\quad e_{z}=\exp(\sum_{\phi\in\Delta^{+}}z_{\phi}{{\mbox{\boldmath{$X$}}}}^{+}_{\phi})e_{0},$
where $e_{0}$ is the extremal weight vector of the representation $\pi$, $\Delta^{+}$ is the set of positive roots of the Lie algebra ${\mathfrak{g}}$, and $X_{\phi}$, $\phi\in\Delta$ $X^{+}_{\phi}$ ($X^{-}_{\phi}$) are the positive (respectively, negative) generators.
In the standard procedure of CS, the $G$-invariant Kähler  two-form on a $2n$-dimensio-
nal homogenous manifold $M=G/H$ is obtained from the Kähler  potential $f$ via the recipe
(3.2a) $-\operatorname{i}\omega_{M}$ =¯$=\partial\bar{\partial}f,~{}f(z,\bar{z})=\log K(z,\bar{z}),~{}K(z,\bar{z}):=(e_{{z}},e_{{z}}),$
(3.2b) $\omega_{M}(z,\bar{z})$ $=\operatorname{i}\sum_{\alpha,\beta}h_{\alpha\bar{\beta}}\operatorname{d}z_{\alpha}\wedge\operatorname{d}\bar{z}_{\beta},~{}h_{\alpha\bar{\beta}}=\frac{\partial^{2}f}{\partial z_{\alpha}\partial\bar{z}_{\beta}},~{}h_{\alpha\bar{\beta}}=\bar{h}_{\beta\bar{\alpha}},~{}\alpha,\beta=1,\dots,n,$
where $K(z,\bar{z})$ is the scalar product of two un-normalized Perelomov’s CS-vectors $e_{{z}}$ at $z\in M$ [18, 26, 79].
It is well known, see [5, Theorem 4.17], [27, Proposition 20], [64, (6), p 156], that the condition
(3.3) $\operatorname{d}\omega=0$
for a Hermitian manifold to have a Kähler  structure is equivalent with the conditions
(3.4) $\frac{\partial h_{\alpha\bar{\beta}}}{\partial z_{\gamma}}=\frac{\partial h_{\gamma\bar{\beta}}}{\partial z_{\alpha}},\quad\text{or}\quad\frac{\partial h_{\alpha\bar{\beta}}}{\partial z_{\gamma}}=\frac{\partial h_{\alpha\bar{\gamma}}}{\partial z_{\bar{\beta}}},\quad\alpha,\beta,\gamma=1,\dots,n.$
In accord with [5, p 42 ], [58, p 28], [27, Appendix B], the Riemannian metric associated with the Hermitian metric on the manifold $M$ in local coordinates is
(3.5) $\operatorname{d}s^{2}_{M}(z,\bar{z})=\sum_{\alpha,\beta}h_{\alpha\bar{\beta}}\operatorname{d}z_{\alpha}\otimes\operatorname{d}\bar{z}_{\beta}.$
Sometimes [48, (7.4)], if the metric is taken as in (3.5), then the Kähler-two form is taken instead of (3.2b) as
(3.6) $-\operatorname{i}\omega_{M}=\frac{\operatorname{i}}{2}\sum_{\alpha,\beta}h_{\alpha\bar{\beta}}\operatorname{d}z_{\alpha}\wedge\operatorname{d}\bar{z}_{\beta}.$
This choice of $f$ in (3.6) corresponds to the situation where the so called $\epsilon$-function [45, 80, 81],
$\epsilon(z):={\mbox{\rm e}}^{-f(z)}K_{M}(z,\bar{z}),$
is constant. The corresponding $G$-invariant metric is called balanced metric. This denomination was firstly used in [54] for compact manifolds, then it was used in [2] for noncompact manifolds, also in [72] in the context of Berezin quantization on homogeneous bounded domain, and we have used it in the case of the partially bounded domain ${\mathcal{D}}^{J}_{n}$ – the Siegel–Jacobi ball [26].
Remark 2.
The Kähler two-form $\omega_{{\mathcal{D}}^{J}_{1}}(w,\eta)$ given by (2.14a) ($\omega(v,\eta)$, (1)) can be obtained from the Kähler  potentials (3.7a) (respectively (3.7b)) using a formula of the type (3.2b)
(3.7a) $f(w,\bar{w},\eta,\bar{\eta})$ $=-2k\log\frac{P}{w}+f(w)+g(\bar{w})+\nu\eta\bar{\eta}+f^{\prime}(\eta)+g^{\prime}(\bar{\eta}),$
(3.7b) $f(v,\bar{v},\eta,\bar{\eta})$ $=-2k\log\frac{v-\bar{v}}{2\operatorname{i}}+f_{1}(v)+g_{1}(\bar{v})+\nu\eta\bar{\eta}+f^{\prime}(\eta)+g^{\prime}(\bar{\eta}).$
3.2. Berry phase on homogenous Kähler  manifolds
Proposition 4.
Let $H$ be the Hamiltonian of a quantum system $(\Psi,{\mathfrak{H}},(,))$ on the homogeneous manifold $M=G/H$ governed by the Schrödinger equation
$H\Psi=\operatorname{i}\dot{\Psi}.$
Let us introduce the notation
(3.8) $\Psi=e^{\operatorname{i}\varphi}\tilde{e}_{z},\quad\varphi\in[0,2\pi).$
Then the phase $\varphi$ is the sum [84]
$\varphi=\varphi_{D}+\varphi_{B}$
of the dynamical $\varphi_{D}$ and the non-adiabatic Berry phase $\varphi_{B}$, where
$\varphi_{D}=-\int{\mathcal{H}}(t)\operatorname{d}t,$
and ${\mathcal{H}}$ is the energy function attached to the Hamiltonian $H$
(3.9) ${\mathcal{H}}=(\tilde{e}_{z}|H|\tilde{e}_{z}).$
The Berry phase is the integral of the one-form $A_{B}$, called Berry connection
(3.10) $\varphi_{B}=\oint A_{B},$
where
(3.11) $\begin{split}A_{B}&\!=\!\frac{\operatorname{i}}{2}\sum_{\alpha\in\Delta_{+n}}(\operatorname{d}z_{\alpha}\partial_{\alpha}\!-\!\operatorname{d}{\bar{z}_{\alpha}}\bar{\partial}_{\alpha})\log(e_{z},e_{z})\!=\!-\operatorname{Im}\theta_{L},\\ \theta_{L}&:\!=\!\sum_{\alpha\in\Delta_{+n}}\partial_{\alpha}f(z,\bar{z})\operatorname{d}{z}_{\alpha}\!=\!\sum_{\alpha\in\Delta_{+n}}\partial_{\alpha}\log(e_{z},e_{z})\operatorname{d}z_{\alpha}\!=\!\sum_{\alpha\in\Delta_{+n}}\frac{\partial_{\alpha}(e_{z},e_{z})}{(e_{z},e_{z})}\operatorname{d}z_{\alpha},\end{split}$
and $f$ is the Kähler  potential defined in (3.2a). The Berry phase depend on the path and not on the Hamiltonian. Closed paths in $M$ imply line integral over connection on the closed paths and are obtained through horizontal lift. If the motion is done on a closed path in M, it generates in the fiber in M the holonomy
(3.12) $\beta=\oint A_{B}=\int_{S}\operatorname{d}A_{B},$
where $\operatorname{d}A_{B}$ is the curvature of the fiber bundle, a realisation of the two-form $V$ of Simon [83],
(3.13) $\begin{split}\operatorname{d}A_{B}&=\frac{\operatorname{i}}{2}\sum_{\alpha,\beta}(-\frac{\partial^{2}f}{\partial{z}_{\beta}\partial\bar{z}_{\alpha}}\operatorname{d}z_{\beta}\wedge\operatorname{d}\bar{z}_{\alpha}+\frac{\partial^{2}f}{\partial\bar{z}_{\beta}\partial z_{\alpha}}\operatorname{d}\bar{z}_{\beta}\wedge\operatorname{d}z_{\alpha})\\ &=-\operatorname{i}\sum_{\alpha,\beta}\frac{\partial^{2}\log(e_{z},e_{z})}{\partial z_{\alpha}\partial\bar{z}_{\beta}}\operatorname{d}z_{\alpha}\wedge\operatorname{d}\bar{z}_{\beta}=-\omega_{M}(z,\bar{z}).\end{split}$
Proof.
The Proposition is taken from [35, (4.17)], [31, Proposition], [21, corrected Proposition 4.1], see also [57, (15)]. The expresion (3.10) of the Berry phase corresponds to the parallel transport, i.e. the vector
(3.14) $|\underline{Z})=e^{\operatorname{i}\varphi_{B}}\tilde{e}_{z},\quad\tilde{e}_{z}:=(e_{z},e_{z})^{-\frac{1}{2}}e_{z}$
in (3.8) has the property that $(\underline{Z},\dot{\underline{Z}})=0$ [31, page 2365] and in the proof it is used the relation
(3.15) $\dot{e}_{z}=\sum_{\alpha}\frac{\partial e_{z}}{\partial z_{\alpha}}\dot{z}_{\alpha}\quad\text{or}~{}~{}\operatorname{d}e_{z}=\partial\operatorname{d}e_{z}=\sum_{\alpha}\frac{\partial e_{z}}{\partial z_{\alpha}}\operatorname{d}z_{\alpha}.$
The proof of the expression (3.13) is a consequence of the relation:
$f=\sum_{j=1}^{n}f_{j}\operatorname{d}x_{j}\Longrightarrow\operatorname{d}f=\sum_{j=1}^{n}\sum_{i=1}^{n}\frac{\partial f_{j}}{\partial x_{i}}\operatorname{d}x_{i}\wedge\operatorname{d}x_{j},$
where $f$ is a smooth function $x_{1},\dots,x_{n}$. The last expression of $\operatorname{d}A_{B}$ in (3.13) is in the convention (3.2b).
Remark 3.
Equation (3.11) of $A_{B}$ can be written with formula (3.15) as
(3.16) $A_{B}=-\operatorname{Im}\frac{(e_{z}|\partial|e_{z})}{(e_{z},e_{z})},~{}~{}\partial f=\frac{\partial f}{\partial z_{\alpha}}\operatorname{d}z_{\alpha}.$
Equation (3.16) is exactly [84, (16) page 10] or [50, (2.56)]
(3.17) $A^{(n)}=-\operatorname{Im}\frac{<n|\partial|n>}{<n|n>},\quad<n|n>=1.$
Equation (3.13) of $\operatorname{d}A_{B}$ can be written as
$\operatorname{d}A^{(n)}=-\operatorname{Im}\frac{(\partial e_{z}|\wedge|\partial e_{z})}{(e_{z},e_{z})}.$
Equation (3.13) of $\operatorname{d}A^{(n)}$ can be written as [84, (13) page 10] or [50, (2.62), (2.63)]
(3.18) $\begin{split}F^{(n)}&=\operatorname{d}A^{(n)}=-\operatorname{Im}\frac{<\operatorname{d}n|\wedge|\operatorname{d}n>}{<n|n>}=\frac{1}{2}\frac{F^{(n)}_{ij}}{<n|n>}\operatorname{d}x_{i}\wedge\operatorname{d}x_{j}\\ &=-\operatorname{Im}(\frac{<\partial_{i}n|\partial_{j}n>-<\partial_{j}n|\partial_{i}n>}{<n|n>})\operatorname{d}x_{i}\wedge\operatorname{d}x_{j},<n|n>=1.\end{split}$
3.3. Linear Hamiltonian in the generators of the Jacobi group $G^{J}_{1}(\mathbb{R})$
The content of the following Remark is mostly extracted from [21, § 4, Lemma 4.1], [30, § 4.3]
Remark 4.
Let us consider a linear Hermitian Hamiltonian $H$ in the generators of the Jacobi group $G^{J}_{1}$
[…]
where
$\bar{\epsilon}_{+}=\epsilon_{-},~{}~{}\epsilon_{a}:=a+\operatorname{i}b,~{}~{}\epsilon_{+}:=m-\operatorname{i}n,~{}~{}\epsilon_{0}:=2c,~{}~{}a,b,c,m,n\in\mathbb{R}.$
The energy function ${\mathcal{H}}$ (3.9) associated to the Hamiltonian (3.19) expressed in the variables $(\eta,v)$ splits into the sum of two independent functions
(3.20) ${\mathcal{H}}(\eta,v)={\mathcal{H}}(\eta)+{\mathcal{H}}(v),\quad v=x+\operatorname{i}y,~{}y>0,~{}\eta=q+\operatorname{i}p,$
where
(3.21a) ${\mathcal{H}}(q,p)$ $=\nu[(m+c)q^{2}+(c-m)p^{2}-2nqp+2(aq+bp)],$
(3.21b) ${\mathcal{H}}(x,y)$ $=\frac{k}{y}[(m+c)(x^{2}+y^{2})+2nx+c-m].$
We particularize equations [30, (3.7)] to the linear Hamiltonian (3.20) to which we add a term ${h}(\kappa)$
(3.22) ${\mathcal{H}}={\mathcal{H}}(p,q)+{\mathcal{H}}(x,y)+{h}(\kappa),$
and we get the equations of motion on the extended Siegel-Jacobi upper half-plane organized as generalized transitive almost cosymplectic manifold $(\tilde{{\mathcal{X}}}^{J}_{1},\theta,\omega)$ corresponding to the energy function (3.22)
(3.23a) $\dot{x}$ $=(c+m)(-x^{2}+y^{2})-2nx-c+m,\quad\dot{y}=-2y[(c+m)x+n],$
(3.23b) $\dot{q}$ $=(c-m)p-qn+b-\frac{q}{2\nu}\frac{\partial h}{\partial\kappa},\quad\dot{p}=\!-\!(m+c)q+\!np-\!a\ \!-\!\frac{p}{2\nu}\frac{\partial{h}}{\partial\kappa},\quad$
(3.23c) $\dot{\kappa}$ $=(c+m)q^{2}+(c-m)p^{2}+aq+bp\!-2npq-\!\frac{1}{\sqrt{\delta}}{\mathcal{H}}$
(3.23d) $=(m+c)[(1-\frac{\nu}{\sqrt{\delta}})q^{2}-\frac{k}{\sqrt{\delta}}\frac{x^{2}+y^{2}}{y}]+(c-m)[(1-\frac{\nu}{\sqrt{\delta}})p^{2}-\frac{k}{\sqrt{\delta}}\frac{1}{y}]$
$-2(1-\frac{\nu}{\sqrt{\delta}})npq+(1-\frac{2}{\sqrt{\delta}})(aq+bp)-2\frac{k}{\sqrt{\delta}}\frac{x}{y}-\frac{1}{\sqrt{\delta}}h.$
Proof.
The differential action of the generators the Jacobi algebra ${\mathfrak{g}}^{J}_{1}$ is given by the formulas:
[…]
(3.24b) ${\mathbb{K}}_{-}=\frac{\partial}{\partial w};~{}{\mathbb{K}}_{0}=k+\frac{1}{2}z\frac{\partial}{\partial z}+w\frac{\partial}{\partial w};$
(3.24c) ${\mathbb{K}}_{+}=\frac{1}{2}\mu z^{2}+2kw+zw\frac{\partial}{\partial z}+w^{2}\frac{\partial}{\partial w},$
where $z\in\mathbb{C}$, $|w|<1$.
Then with (2.9), we get
(3.25) $\frac{\partial K}{\partial z}=\nu\bar{\eta}K,\quad\frac{\partial K}{\partial w}=(2k\frac{\bar{w}}{1-w\bar{w}}+\frac{\nu}{2}\bar{\eta}^{2})K,$
and $\eta$ was defined in (3.9).
With (4) we find
(3.26a) $\frac{\partial H(x,y)}{\partial x}=$ $\frac{2k}{y}[(m\!+\!c)x\!+\!n];\frac{\partial H(x,y)}{\partial y}=\!\frac{k}{y^{2}}[(m\!+\!c)(y^{2}-x^{2})\!+\!m\!-\!c\!-\!2nx];$
(3.26b) $\frac{\partial H(q,p)}{\partial q}=$ $2\nu[(m\!+\!c)q\!-\!np\!+\!a];\frac{\partial H(q,p)}{\partial p}\!=\!2\nu[(c\!-\!m)p\!-\!nq\!+\!b].$
[30, (4.5)-(4.7)] and (3.3) imply
(3.27a) $\dot{q}_{1}$ $=k\dot{x}=A_{1}=y^{2}\frac{\partial H}{\partial y}=k[(m+c)(y^{2}-x^{2})+m-c-2nx]$
(3.27b) $\dot{q}_{2}$ $=2\nu\dot{q}=A_{2}=\frac{\partial H}{\partial p}-\frac{q}{2\nu}\frac{\partial H}{\partial\kappa}=2\nu[{(c-m)p-nq+b}-q]-q\frac{\partial H}{\partial\kappa}$
(3.27c) $\dot{p}_{1}$ $=y^{-2}\dot{y}=B_{1}=-\frac{1}{k}\frac{\partial H}{\partial x}=-\frac{1}{y}[(c+m)x+n]$
(3.27d) $\dot{p}_{2}$ $=\dot{p}=B_{2}=-\frac{\nu}{2\nu}[\frac{\partial H}{\partial q}+p\frac{\partial H}{\partial k}]=-\frac{2}{2\nu}[(c+m)q-np+a]-\frac{1}{2\nu}p\frac{\partial H}{\partial k}$
(3.27e) $\dot{\kappa}$ $=\frac{1}{2\nu}(p\frac{\partial H}{\partial p}+q\frac{\partial H}{\partial q})-\frac{1}{\sqrt{\delta}}H,$
and (4) are proved. ∎
3.4. Kähler two-forms, Christofell’s symbols and connections matrices on ${\mathcal{D}}_{1}$ and ${\mathcal{X}}_{1}$
a) The Kähler two form on the Siegel disk ${\mathcal{D}}_{1}$ corresponding to the Kähler  potential
$f(w)=-2k\log P$
is [15, (7.21)]
(3.28) $-\operatorname{i}\omega(w,\bar{w})=\frac{2k}{P^{2}}\operatorname{d}w\wedge\operatorname{d}\bar{w}.$
If we make the change of variables (2.28b) $w\rightarrow v$ and apply
(3.29) $\operatorname{d}w=2\operatorname{i}\frac{\operatorname{d}v}{(v+i)^{2}},$
we get
(3.30a) $-\operatorname{i}\omega(v,\bar{v})$ $=-\frac{2k}{(v-\bar{v})^{2}}\operatorname{d}v\wedge\operatorname{d}\bar{v},=\frac{k}{2y^{2}}\operatorname{d}v\wedge\operatorname{d}\bar{v},$
(3.30b) $\omega(x,y)$ $=\frac{k}{y^{2}}\operatorname{d}x\wedge\operatorname{d}y,~{}~{}v=x=\operatorname{i}y.$
Alternatively, if in (3.28) we introduce (2.15) and then (3.7), we get again (3.4).
b) From (2.11) we get
$h_{{\mathcal{D}}^{J}_{1}}(w)=\frac{2k}{P^{2}}.$
With formula (5.39) we get
(3.31) $\Gamma^{w}_{ww}=\frac{2\bar{w}}{P}$
We consider formula (1) of the transformation $w\rightarrow v$ and we write this change of variables as
(3.32) $\Gamma^{v}_{vv}=\Gamma^{w}_{ww}\frac{\partial w}{\partial v}+\frac{\partial^{2}w}{\partial v^{2}}\frac{\partial v}{\partial w}.$
With (3.29) we get $\frac{\partial w}{\partial v}$ and then
$\frac{\partial^{2}w}{\partial v^{2}}=-\frac{4\operatorname{i}}{(v+\operatorname{i})^{3}}.$
We get with (3.32)
(3.33) $\Gamma^{v}_{vv}=\frac{2}{\bar{v}-v}=\frac{\operatorname{i}}{y},$
which is correct, because if we apply (5.39) to
(3.34) $h_{{\mathcal{X}}^{J}_{1}}(v)=-\frac{2k}{(v-\bar{v})^{2}}=\frac{k}{2y^{2}},$
we get (3.33).
We have
(3.35) $v-\bar{v}=2\operatorname{i}\frac{P}{(1-w)(1-\bar{w})}.$
Inverse, if we consider the change of variables $v\rightarrow w$ and apply the inverse formula to (3.32)
$\Gamma^{w}_{ww}=\Gamma^{v}_{vv}\frac{\partial v}{\partial w}+\frac{\partial^{2}v}{\partial w^{2}}\frac{\partial w}{\partial v}.$
starting with (3.33) we get (3.31).
c) With formula (3.33) we get
(3.36) $\theta^{v}_{v}(v)=\Gamma^{v}_{vv}\operatorname{d}v=\frac{2}{v-\bar{v}}\operatorname{d}v.$
We apply relation (5.7) to the change of variables $v\rightarrow w$ (1). We have
$A(w)=\frac{2\operatorname{i}}{(1-w)^{2}},\quad\operatorname{d}AA^{-1}=2\frac{\operatorname{d}w}{1-w}$
and we should have
(3.37) $\omega^{\prime}=\omega_{J^{1}}(w)=2\frac{\operatorname{d}w}{1-w}+\frac{2}{v-\bar{v}}\frac{2\operatorname{i}}{(1-w)^{2}}\operatorname{d}w.$
With (3.35) we find
$\theta^{w}_{ww}=\Gamma^{w}_{ww}\operatorname{d}w=2\frac{\bar{w}}{1-w\bar{w}}\operatorname{d}w,$
in accord with (3.31).
3.5. Connection matrices on ${\mathcal{D}}^{J}_{1}$ and ${\mathcal{X}}^{J}_{1}$
Christoffel’s symbols $\Gamma$-s have the expressions [25, (38)]
(3.38) $\begin{split}\Gamma^{z}_{zz}&=-\lambda\bar{\eta};~{}\Gamma^{w}_{zz}=\lambda;~{}\Gamma^{z}_{zw}=-\lambda\bar{\eta}^{2}+\frac{\bar{w}}{P};\\ \Gamma^{w}_{wz}&=\lambda\bar{\eta};~{}\Gamma^{z}_{ww}=-\lambda\bar{\eta}^{3};~{}\Gamma^{w}_{ww}=\lambda\bar{\eta}^{2}+2\frac{\bar{w}}{P},~{}\lambda=\frac{\nu}{2k}.\end{split}$
The connection matrix (form) $\theta_{{\mathcal{D}}^{J}_{1}}$ on ${\mathcal{D}}^{J}_{1}$ in $(w,z)$ is [25, (40)]
(3.39e) $\theta_{{\mathcal{D}}^{J}_{1}}(w,z):$ $=\left(\begin{array}[]{cc}\theta^{w}_{w}&\theta^{z}_{w}\\ \theta^{w}_{z}&\theta^{z}_{z}\end{array}\right)=\left(\begin{array}[]{cc}\Gamma^{w}_{wz}\operatorname{d}z+\Gamma^{w}_{ww}\operatorname{d}w&\Gamma^{z}_{wz}\operatorname{d}z+\Gamma^{z}_{ww}\operatorname{d}w\\ \Gamma^{w}_{zz}\operatorname{d}z+\Gamma^{w}_{zw}\operatorname{d}w&\Gamma^{z}_{zz}\operatorname{d}z+\Gamma^{z}_{zw}\operatorname{d}w\end{array}\right)$
(3.39h) $=\left(\begin{array}[]{cc}\lambda\bar{\eta}{\mathcal{A}}+2\frac{\bar{w}}{P}\operatorname{d}w&-\lambda\bar{\eta}^{2}{\mathcal{A}}+\frac{\bar{w}}{P}\operatorname{d}z\\ \lambda{\mathcal{A}}&-\lambda\bar{\eta}{\mathcal{A}}+\frac{\bar{w}}{P}\operatorname{d}w\end{array}\right).$
In the variables $(u,v)\in(\mathbb{C},{\mathcal{X}}_{1})$ the geodesic equations (5.43) for the metric (2.35) read [29, (57)]
(3.40) $\left\{\begin{array}[]{l}\frac{\operatorname{d}^{2}u}{\operatorname{d}t^{2}}+\Gamma^{u}_{uu}\left(\frac{\operatorname{d}u}{\operatorname{d}t}\right)^{2}+2\Gamma^{u}_{uv}\frac{\operatorname{d}u}{\operatorname{d}t}\frac{\operatorname{d}v}{\operatorname{d}t}+\Gamma^{u}_{vv}\left(\frac{\operatorname{d}v}{\operatorname{d}t}\right)^{2}=0;\\ \frac{\operatorname{d}^{2}v}{\operatorname{d}t^{2}}+\Gamma^{v}_{uu}\left(\frac{\operatorname{d}u}{\operatorname{d}t}\right)^{2}+2\Gamma^{v}_{uv}\frac{\operatorname{d}u}{\operatorname{d}t}\frac{\operatorname{d}v}{\operatorname{d}t}+\Gamma^{v}_{vv}\left(\frac{\operatorname{d}v}{\operatorname{d}t}\right)^{2}=0.\end{array}\right.$
Christoffel’s symbols $\Gamma$-s in $(u,v)$ corresponding to the Riemannian metric associated to the Kähler two-form (2.31a) are extracted from [29, (62)] with corrections
(3.41) $\begin{split}\Gamma^{u}_{uu}&=\frac{\operatorname{i}}{\iota}r,~{}\Gamma^{v}_{uu}=\frac{\operatorname{i}}{\iota},~{}\Gamma^{u}_{uv}=\frac{\operatorname{i}}{2\iota}(\frac{\iota}{y}-2r^{2});\\ \Gamma^{v}_{vu}&=-\frac{\operatorname{i}}{\iota}r,~{}\Gamma^{u}_{vv}=\frac{\operatorname{i}}{\iota}r^{3},~{}\Gamma^{v}_{vv}=\frac{\operatorname{i}}{\iota}(\frac{\iota}{y}+r^{2}),\end{split}$
where
(3.42) $\iota=\frac{k}{\nu}=\frac{1}{2\lambda},\quad r=\frac{n}{y}.$
Equations (3.40) with the $\Gamma$-s (3.41) lead to the same equations as [29, (53)]
(3.43a) $\ddot{v}+\frac{\operatorname{i}}{\iota}\left[\dot{u}^{2}-2r\dot{u}\dot{v}+(\frac{\iota}{y}+r^{2})\dot{v}^{2}\right]=0,$
(3.43b) $\ddot{u}+\frac{\operatorname{i}}{\iota}\left[r\dot{u}^{2}+(\frac{\iota}{y}-2r^{2})\dot{u}\dot{v}+r^{3}\dot{v}^{2}\right]=0.$
We check the value of $\Gamma^{u}_{uv}$. We have
$\left\{\begin{array}[]{cc}h_{v\bar{u}}\Gamma^{v}_{vu}+h_{u\bar{u}}\Gamma^{u}_{vu}&=\frac{\partial h_{u\bar{u}}}{\partial v}\\ h_{v\bar{v}}\Gamma^{v}_{vu}+h_{u\bar{v}}\Gamma^{u}_{vu}&=\frac{\partial h_{u\bar{v}}}{\partial v}\end{array}\right..$
$\left\{\begin{array}[]{cc}-\frac{\nu r}{y}\Gamma^{u}_{vu}+\frac{\nu}{y}\Gamma^{u}_{vu}&=\frac{\operatorname{i}\nu}{2y^{2}}\\ (\frac{k}{2y^{2}}+\frac{\nu r^{2}}{y})\Gamma^{v}_{vu}-\frac{\nu r}{y}\Gamma^{u}_{vu}&=-\frac{\operatorname{i}\nu r}{y^{2}}\end{array}\right..$
With the notation
$\Delta=-\left\|\begin{array}[]{cc}h_{u\bar{u}}&h_{v\bar{u}}\\ h_{u\bar{v}}&h_{v\bar{v}}\end{array}\right\|=-\frac{\nu k}{2y^{3}};$
$\Delta_{1}=\left\|\begin{array}[]{cc}\frac{\operatorname{i}\nu}{2y^{2}}&\frac{\nu}{y}\\ -\frac{\operatorname{i}\nu r}{y^{2}}&-\frac{\nu r}{y}\end{array}\right\|=\frac{\operatorname{i}\nu^{2}r}{2y^{3}};$
$\Delta_{2}=-\left\|\begin{array}[]{cc}\frac{\operatorname{i}\nu}{2y^{2}}&-\frac{\nu r}{y}\\ -\frac{\operatorname{i}\nu r}{y^{2}}&\frac{k}{2y^{2}}+\frac{\nu r^{2}}{y}\end{array}\right\|=-\operatorname{i}\frac{\nu k}{4y^{4}}+\frac{\operatorname{i}\nu^{2}r^{2}}{2y^{3}},$
we get
$\Gamma^{v}_{vu}=\frac{\Delta_{1}}{\Delta}=-\frac{\nu r}{k};~{}~{}~{}\Gamma^{u}_{vu}=\frac{\operatorname{i}}{2y}-\frac{\operatorname{i}}{\iota}r^{2}.$
We also have
$\Gamma^{u}_{vu}=h^{v\bar{u}}\frac{\partial h_{u\bar{v}}}{\partial v}+h^{u\bar{u}}\frac{\partial h_{u\bar{u}}}{\partial v}=\frac{2y^{2}r}{k}\frac{\partial}{\partial v}(-\frac{\nu r}{y})+[\frac{y}{\nu}+\frac{\nu r^{2}y^{2}}{k}\frac{\partial}{\partial v}(\frac{\nu}{y})]=\frac{1}{2}\frac{\operatorname{i}}{\iota}(\frac{\iota}{y}-2r^{2}).$
We get for the connection matrix on ${\mathcal{X}}^{J}_{1}$ in the variables $(v,u)$ the expression
(3.44e) $\theta_{{\mathcal{X}}^{J}_{1}}(v,u)$ $=\left(\begin{array}[]{cc}\theta^{v}_{v}&\theta^{u}_{v}\\ \theta^{v}_{u}&\theta^{u}_{u}\end{array}\right)=\left(\begin{array}[]{cc}\Gamma^{v}_{vu}\operatorname{d}u+\Gamma^{v}_{vv}\operatorname{d}v&\Gamma^{u}_{vu}\operatorname{d}u+\Gamma^{u}_{vv}\operatorname{d}v\\ \Gamma^{v}_{uu}\operatorname{d}u+\Gamma^{v}_{uv}\operatorname{d}v&\Gamma^{u}_{uu}\operatorname{d}u+\Gamma^{u}_{uv}\operatorname{d}v\end{array}\right)$
(3.44h) $\!=\!\frac{\operatorname{i}}{\iota}\left(\begin{array}[]{cc}-r{\mathcal{B}}+\frac{\iota}{y}\operatorname{d}v&-r^{2}{\mathcal{B}}+\frac{\iota}{2y}\\ {\mathcal{B}}&r{\mathcal{B}}+\frac{\iota}{2y}\operatorname{d}v\end{array}\right)\!$
(3.44m) $=\!\frac{\operatorname{i}}{\iota}\!\left[\left(\begin{array}[]{cc}-r&-r^{2}\\ 1&r\end{array}\right){\mathcal{B}}\!+\!\frac{\iota}{2y}\left(\begin{array}[]{cc}2\operatorname{d}v&\operatorname{d}u\\ 0&\operatorname{d}v\end{array}\right)\!\right].$
Now we calculate the connection matrix on the Siegel-Jacobi disc in the variables $(x,y,q,p)$.
The non-zero Christoffel’s symbols corresponding to the Riemannian metric $\operatorname{d}\!s^{2}_{{\mathcal{X}}^{J}_{1}}\!(\!x\!,y\!,\!q,\!p\!)$ (1) on the Siegel-Jacobi upper half-plane are [29, (73)]
(3.45) $\begin{array}[]{lllll}\Gamma^{x}_{xy}=-\frac{1}{y}&\Gamma^{x}_{pp}=-\epsilon\,xy&\Gamma^{x}_{pq}=-\frac{1}{2}\epsilon\,y&&\\ \Gamma^{y}_{xx}=\frac{1}{y}&\Gamma^{y}_{yy}=-\frac{1}{y}&\Gamma^{y}_{pp}=\frac{\epsilon\,}{2}(x^{2}\!\!-\!\!y^{2})&\Gamma^{y}_{pq}=\frac{\epsilon\,}{2}x&\Gamma^{y}_{qq}=\frac{\epsilon\,}{2}\\ \Gamma^{p}_{xp}=\frac{1}{2}\frac{x}{y^{2}}&\Gamma^{p}_{xq}=\frac{1}{2}\frac{1}{y^{2}}&\Gamma^{p}_{yp}=\frac{1}{2y}&&\\ \Gamma^{q}_{xp}=\frac{y^{2}-x^{2}}{2y^{2}}&\Gamma^{q}_{xq}=-\frac{x}{2y^{2}}&\Gamma^{q}_{yp}=-\frac{x}{y}&\Gamma^{q}_{yq}=-\frac{1}{2y}&\end{array},$
where
(3.46) $\epsilon=\frac{\gamma}{\alpha}=2\frac{\nu}{k}.$
Let
(3.47) $\!\theta_{{\mathcal{X}}^{J}_{1}}(x,y,q,p)\!=\!\left(\begin{array}[]{cccc}\theta^{x}_{x}&\theta^{x}_{y}&\theta^{x}_{q}&\theta^{x}_{p}\\ \theta^{y}_{x}&\theta^{y}_{y}&\theta^{y}_{q}&\theta^{y}_{p}\\ \theta^{q}_{x}&\theta^{q}_{y}&\theta^{q}_{q}&\theta^{q}_{p}\\ \theta^{p}_{x}&\theta^{p}_{y}&\theta^{p}_{q}&\theta^{p}_{p}\end{array}\right)\!.$
We find for the matrix elements of (3.47) the values
(3.48e) $\theta_{{\mathcal{X}}^{J}_{1}}=$ $\left(\begin{array}[]{cccc}\!-\!\frac{\operatorname{d}y}{y}&\!-\!\frac{\operatorname{d}x}{y}&\!-\frac{\epsilon\,}{2}y\operatorname{d}p&\!-\epsilon\,xy\operatorname{d}p\!-\!\frac{\epsilon\,y}{2}\operatorname{d}q\\ \frac{\operatorname{d}x}{y}&\!-\!\frac{\operatorname{d}y}{y}&\frac{\epsilon\,}{2}\operatorname{d}q\!+\!\frac{\epsilon\,}{2}x\operatorname{d}p&\!\frac{\epsilon\,}{2}x\operatorname{d}q\!+\!\frac{\epsilon\,}{2}(x^{2}-y^{2})\operatorname{d}p\\ \!-\!\frac{x}{2y^{2}}\operatorname{d}q\!+\!\frac{y^{2}\!-\!x^{2}}{2y^{2}}\operatorname{d}p&\!-\!\frac{x}{y}\operatorname{d}y&\!-\!\frac{x}{2y^{2}}\operatorname{d}x&\frac{y^{2}\!-\!x^{2}}{2y^{2}}\operatorname{d}x\!-\!\frac{x}{y}\operatorname{d}y\\ \frac{x}{2y^{2}}\operatorname{d}p\!+\!\frac{1}{2y^{2}}\operatorname{d}q&\frac{1}{2y}\operatorname{d}p&\frac{1}{2y^{2}}\operatorname{d}x&\frac{x}{2y^{2}}\operatorname{d}x\!+\!\frac{1}{2y}\operatorname{d}y\end{array}\right)$
(3.48n) $=$ $\left(\begin{array}[]{cccc}(\!0\!,-\frac{1}{y},\!0\!,\!0)&(-\frac{1}{y},\!0\!,\!0,\!0)&(\!0,\!0,\!0,\!-\!\frac{\epsilon\,y}{2})&(\!0\!,\!0\!,-\!\frac{\epsilon\,}{2}y,-\epsilon\,xy)\\ (\frac{1}{y}\!,\!0\!,\!0\!,\!0)&(\!0\!,\!-\frac{1}{y}\!,\!0\!,\!0\!)&(\!0\!,\!0\!,\!\frac{\epsilon\,}{2}\!,\!\frac{\epsilon\,x}{2})&(\!0\!,0\!,\frac{\epsilon\,x}{2}\!,\!\frac{\epsilon\,}{2}(x^{2}\!-\!y^{2}))\\ (\!0\!,\!0\!,-\frac{x}{2y^{2}}\!,\!\frac{y^{2}\!-\!x^{2}}{2y^{2}})&(\!0\!,-\frac{x}{y},\!0\!,\!0\!)&(\!-\frac{x}{2y^{2}},\!0\!,\!0\!,\!0\!)&(\frac{y^{2}\!-\!x^{2}}{2y^{2}},-\frac{x}{y},\!0\!,\!0)\\ (\!0,\!0,\frac{1}{2y^{2}}\!,\!\frac{x}{2y^{2}})&(\!0\!,\!0\!,\!0\!,\!\frac{1}{2y})&(\frac{1}{2y^{2}},\!0\!,\!0\!,\!0)&(\frac{x}{2y^{2}},\frac{1}{2y},\!0\!,\!0\!)\end{array}\right)\!\otimes\!\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\end{array}\right)$
3.6. Check of formulae (5.23a)
We write (5.23a) as
$\omega^{\prime}(v,u)=\operatorname{d}J_{WU}J^{-1}_{WU}+J_{WU}\omega(w,z)J^{-1}_{WU},$
where $(w,z)$ ($(v,u)$) are the “old” coordinates (respectively “new” coordinates). If denote $J_{WU}$ with $A=A((w,z)\rightarrow(v,u))$, we have
$\left(\begin{array}[]{c}\frac{\partial}{\partial v}\\ \frac{\partial}{\partial u}\end{array}\right)=A\left(\begin{array}[]{c}\frac{\partial}{\partial w}\\ \frac{\partial}{\partial z}\end{array}\right)=\left(\begin{array}[]{cc}\frac{\partial w}{\partial v}&\frac{\partial z}{\partial v}\\ \frac{\partial w}{\partial u}&\frac{\partial z}{\partial u}\end{array}\right)\left(\begin{array}[]{c}\frac{\partial}{\partial w}\\ \frac{\partial}{\partial z}\end{array}\right)$
For the partial Cayley transform (2.28a) we find
(3.49) $A=\left(\begin{array}[]{cc}\frac{2\operatorname{i}}{(v+\operatorname{i})^{2}}&-\frac{2\operatorname{i}u}{(v+\operatorname{i})^{2}}\\ 0&\frac{2\operatorname{i}}{v+\operatorname{i}}\end{array}\right);~{}~{}~{}A^{-1}=-\frac{\operatorname{i}}{2}(v+\operatorname{i})\left(\begin{array}[]{cc}v+\operatorname{i}&u\\ 0&1\end{array}\right).$
(3.50) $\operatorname{d}A=\frac{2\operatorname{i}}{(v+\operatorname{i})^{2}}\left(\begin{array}[]{cc}-\frac{2\operatorname{d}v}{v+\operatorname{i}}&\frac{-(v+\operatorname{i})\operatorname{d}u+2u\operatorname{d}v}{v+\operatorname{i}}\\ 0&-\operatorname{d}v\end{array}\right).$
(3.51) $\operatorname{d}A\cdot A^{-1}=-\frac{1}{v+\operatorname{i}}\left(\begin{array}[]{cc}2\operatorname{d}v&\operatorname{d}u\\ 0&\operatorname{d}v\end{array}\right)$
(3.52) $J_{UW}\omega(w,z)J^{-1}_{UW}\!=\!\left(\begin{array}[]{cc}\lambda{\mathcal{A}}(\bar{\eta}-u)\!+\!2\frac{\bar{w}}{P}\operatorname{d}w&\frac{-\lambda{\mathcal{A}}(u-\bar{\eta})^{2}\!+\!\frac{\bar{w}}{P}(u\operatorname{d}w+\operatorname{d}z)}{v+\operatorname{i}}\\ \lambda{\mathcal{A}}(v+\operatorname{i})&\lambda{\mathcal{A}}(u-\bar{\eta})\!+\!\frac{\bar{w}}{P}\operatorname{d}w\end{array}\right).$
(3.53) $\omega^{\prime}(v,u)=\!\left(\begin{array}[]{cc}\lambda{\mathcal{A}}(\bar{\eta}-u)\!+\!2\frac{\bar{w}}{P}\operatorname{d}w-\frac{2\operatorname{d}v}{v+\operatorname{i}}&\frac{-\lambda{\mathcal{A}}(u-\bar{\eta})^{2}\!+\!\frac{\bar{w}}{P}(u\operatorname{d}w+\operatorname{d}z)}{v+\operatorname{i}}-\frac{\operatorname{d}u}{v+\operatorname{i}}\\ \lambda{\mathcal{A}}(v+\operatorname{i})&\lambda{\mathcal{A}}(u-\bar{\eta})\!+\!\frac{\bar{w}}{P}\operatorname{d}w-\frac{\operatorname{d}v}{v+\operatorname{i}}\end{array}\right).$
Now we compare equations (3.5), (3.53).
We compare firstly the terms “21”. Because of (2.30), (3.42), it is verified that
$\frac{\operatorname{i}}{\iota}{\mathcal{B}}=\lambda{\mathcal{A}}(v+\operatorname{i}),$
We compare the terms “11”. We should have
$2\frac{\bar{w}}{P}\operatorname{d}w-2\frac{\operatorname{d}v}{v+\operatorname{i}}=\operatorname{i}\frac{\operatorname{d}v}{y}$
which is true because of (2.28a) and (3.35).
We compare the terms “12”. We should have
(3.54) $\frac{\operatorname{i}}{2y}\operatorname{d}u=-\frac{\operatorname{d}u}{v+\operatorname{i}}+\frac{\bar{w}}{P}\frac{u\operatorname{d}w+\operatorname{d}z}{v+\operatorname{i}},$
which is true.
Indeed, from (2.28a) we have
$\operatorname{d}u=\frac{(1-w)\operatorname{d}z+z\operatorname{d}w}{(1-w)^{2}},\quad v+\operatorname{i}=\frac{2\operatorname{i}}{1-w},$
and also we use
$y=\frac{1-w\bar{w}}{(1-w)(1-bar{w})}.$
We compare the terms “22” . We should have
$\frac{\bar{w}}{P}\operatorname{d}w-\frac{dv}{v+\operatorname{i}}=\operatorname{i}\frac{\operatorname{d}v}{2y},$
$\frac{\bar{w}}{P}\operatorname{d}w=(\frac{\operatorname{i}}{2y}+\frac{1}{v+\operatorname{i}})\operatorname{d}v,$
which is true because of of (1), (3.35).
3.7. Berry phase on ${\mathcal{D}}^{J}_{1}$ and ${\mathcal{X}}^{J}_{1}$
a) Firstly we calculate the Berry phase on ${\mathcal{X}}_{1}$ from the Berry phase on ${\mathcal{D}}_{1}$.
The relations(2.28b), (2.15), (2.3) and (2.26) implies
(3.55) $\alpha=\frac{x^{2}+y^{2}-1}{N},\quad\beta=-2\frac{x}{N},\quad$
With
$f=-2k\log P=-2k\log(1-|w|^{2}),$
we get with (3.11) the Berry phase on the Sigel disk ${\mathcal{D}}_{1}$
(3.56) $A_{B}(w,\bar{w})=\operatorname{i}k\frac{\bar{w}\operatorname{d}w-w\operatorname{d}\bar{w}}{P}=2k\frac{\beta\operatorname{d}\alpha-\alpha\operatorname{d}\beta}{P}.$
But from (3.55)
(3.57a) $\operatorname{d}\alpha$ $=2\frac{2x(y+1)\operatorname{d}x+[-x^{2}+(y+1)^{2}]\operatorname{d}y}{N^{2}},$
(3.57b) $\operatorname{d}\beta$ $=-2\frac{[-x^{2}+(y+1)^{2}]\operatorname{d}x-2x(y+1)\operatorname{d}y}{N^{2}},$
and
$-\alpha\operatorname{d}\beta+\beta\operatorname{d}\alpha=$ $\frac{2}{N^{3}}\{[(x^{2}+y^{2}-1)(-x^{2}+(y+1)^{2})-4x^{2}(y+1)]\operatorname{d}x$
$+[-2x(x^{2}+y^{2}-1)(y+1)-2x(-x^{2}+(y+1)^{2}]\operatorname{d}y\}.$
(2.12) implies for the Berry phase $\phi_{B}(x,y)$
$\phi_{B}(x,y)=\frac{k}{2}\frac{2}{N^{2}}\frac{-N(x^{2}-y^{2}+1)\operatorname{d}x-2Nxy\operatorname{d}y}{y},$
and Berry phase on ${\mathcal{X}}_{1}$ in $(x,y)$ obtained from the Berry phase on ${\mathcal{D}}_{1}$ in $(\alpha,\beta)$ is
(3.59) $\phi_{B}(x,y)=k\frac{(-x^{2}+y^{2}-1)\operatorname{d}x-2xy\operatorname{d}y}{y[x^{2}+(y+1)^{2}]}.$
Christofell’s symbols in the variables $(x,y)$ on ${\mathcal{X}}_{1}$ are extracted from [29, (69)]
(3.60a) $\Gamma^{x}_{xx}$ $=0,\quad\Gamma^{x}_{xy}=-\frac{1}{y},\quad\Gamma^{x}_{yy}=0,$
(3.60b) $\Gamma^{y}_{xx}$ $=\frac{1}{y},\quad\Gamma^{y}_{xy}=0,\quad\Gamma^{y}_{yy}=-\frac{1}{y}.$
b) The Berry phase on the Siegel-Jacobi disk ${\mathcal{D}}^{J}_{1}$ in $(w,z)$, $(\alpha,\beta,q,p)$.
The starting point is the scalar product of two CS on ${\mathcal{D}}^{J}_{1}$ [15, (7.13b)]
(3.61) $f(z,w)=(e_{z,w},e_{z,w})=-2k\log P+\nu F,\quad F=\frac{2z\bar{z}+\bar{w}z^{2}+w\bar{z}^{2}}{2P}.$
With (3.11) we get
(3.62) $A_{B}(z,w)=\frac{\operatorname{i}}{2}(A(z,w)-cc),$
where [21, §4.2]
(3.63a) $A(z,w)$ $=(2k\frac{\bar{w}}{P}+\frac{\nu}{2}\bar{\eta}^{2})\operatorname{d}w+\nu\bar{\eta}\operatorname{d}z$
(3.63b) $=k(\Gamma^{w}_{ww}\operatorname{d}w+2\Gamma^{w}_{wz}\operatorname{d}z).$
With (3.39e) we rewrite (3.63b) as
$A(z,w)=k(\theta^{w}_{w}+\Gamma^{w}_{wz}\operatorname{d}z),$
which is in fact formula before (4.27) in [21].
With (2.12) we get for $A$ in (3.62)
$A(w,\eta)=(2k\frac{\bar{w}}{P}-\frac{\nu}{2}\bar{\eta}^{2})\operatorname{d}w+\nu\bar{\eta}(\operatorname{d}\eta-w\operatorname{d}\bar{\eta}).$
We also have the following expression for the Berry phase
(3.64) $\begin{split}A_{B}(\alpha,\beta,q,p)&=2k\frac{\beta}{1-\alpha^{2}-\beta^{2}}\operatorname{d}\alpha+[-2k\frac{\alpha}{1-\alpha^{2}-\beta^{2}}+\frac{\nu}{2}(q^{2}-p^{2})]\operatorname{d}\beta\\ &-\nu[(\alpha-\beta q-1)\operatorname{d}q+(\alpha q+\beta)\operatorname{d}p].\end{split}$
c) Berry phase in $(u,v)$
We use for $f(u,v)$ (2.44a), $(u,v)=(m+\operatorname{i}n,x+\operatorname{i}y)$. We have
$f_{v}=-\frac{k}{2}\frac{1}{v-\bar{v}}+\operatorname{i}\nu r^{2}=\operatorname{i}(\frac{k}{4y}+\nu\frac{n^{2}}{y^{2}}),\quad f_{u}=-2\operatorname{i}\nu r=-2\operatorname{i}\nu\frac{n}{y}.$
We get
(3.65a) $A(u,v)$ $=f_{u}\operatorname{d}u+f_{v}\operatorname{d}v=-2\operatorname{i}\nu r\operatorname{d}u+\operatorname{i}(\frac{k}{4y}+\nu r^{2})\operatorname{d}v$
(3.65b) $=k[-2\Gamma^{u}_{uu}\operatorname{d}u+(\Gamma^{v}_{vv}-\frac{3i}{4}\frac{1}{y})\operatorname{d}v],$
where we have used (3.41).
(3.66) $A_{B}(x,y,m,n)=\frac{\operatorname{i}}{2}(A(u,v)-\bar{A}(u,v))=\frac{1}{y}\left[-(\frac{k}{4}+\nu\frac{n^{2}}{y})\operatorname{d}x+2\nu n\operatorname{d}m\right].$
d) Berry phase on ${\mathcal{X}}^{J}_{1}$ in $(v,\eta)=(x+\operatorname{i}y,q+\operatorname{i}p)$
With (4.17) for $f(v,\eta)$, we have
$f_{v}=-2k\frac{1}{v-\bar{v}}=\operatorname{i}\frac{k}{y};\quad f_{\eta}=-\nu(\eta-\bar{\eta})=-2\operatorname{i}\nu p,$
and we get
$A_{B}(x,y,q,p)=-\frac{k}{y}\operatorname{d}x+2\nu p\operatorname{d}q.$
We write
(3.67) $A_{B}(\alpha,\beta,x,y)=\operatorname{d}\phi_{B}(\alpha,\beta,x,y)=\operatorname{d}\phi_{BI}+\operatorname{d}\phi_{BII}+\operatorname{d}\phi_{BIII},$
where $\operatorname{d}\phi_{BI}$, which appears in (3.56), was calculated as (3.59), and
(3.68a) $\operatorname{d}\phi_{BII}$ $=\frac{\operatorname{i}\nu}{2}(-\bar{\eta}^{2}\operatorname{d}w+cc),$
(3.68b) $\operatorname{d}\phi_{BIII}$ $=\frac{\operatorname{i}\nu}{2}[(\bar{\eta}+w\eta)\operatorname{d}\eta-cc].$
We find
(3.69a) $\operatorname{d}\phi_{BII}$ $=\frac{\operatorname{i}\nu}{2}[-(q-\operatorname{i}p)^{2}(\operatorname{d}\alpha+\operatorname{i}\operatorname{d}\beta)+cc]=\nu[(q^{2}-p^{2})\operatorname{d}\beta-2qp\operatorname{d}\alpha],$
(3.69b) $\operatorname{d}\phi_{BIII}$ $=\nu\{-[(\alpha+1)q+\beta p]\operatorname{d}p+[(1-\alpha)p+\beta q]\operatorname{d}q\}.$
With (3.55), (3.7) we get
(3.70) $\begin{split}A_{B}&(x,y,p,q)=\frac{k}{Ny}[(-x^{2}+y^{2}-1)\operatorname{d}x-2xy\operatorname{d}y]+\\ &\frac{2\nu}{N^{2}}[-4x(y+1)pq+(x^{2}-(y+1)^{2})(q^{2}-p^{2})]\operatorname{d}x+\\ &\frac{4\nu}{N^{2}}[(x^{2}-(y+1)^{2})qp+x(y+1)(q^{2}-p^{2})]\operatorname{d}y+\\ &\frac{2\nu}{N}\{-[(x^{2}+y(y+1))q-xp]\operatorname{d}p+[(y+1)p-xq]\operatorname{d}q]\}=\\ &\frac{1}{N}\{\frac{k}{y}(-x^{2}+y^{2}-1)+\frac{2\nu}{N}[-4x(y+1)pq+(x^{2}-(y+1)^{2})(q^{2}-p^{2})]\}\operatorname{d}x+\\ &\frac{2}{N}\{-kx+\frac{2\nu}{N}[x^{2}-(y+1)^{2}]pq+x(y+1)(q^{2}-p^{2})\}\operatorname{d}y+\\ &\frac{2\nu}{N}\{[-[x^{2}+y(y+1)]q+xp]\operatorname{d}p+[(y+1)p-xq]\operatorname{d}q\}.\end{split}$
4. Almost cosymplectic manifolds
4.1. Definitions
Following [70], an almost cosymplectic manifold (ACOS) is the triplet $(M,\theta,\Omega)$, where $M$ is a $(2{\mathrm{n}}+1)$-dimensional manifold, $\theta\in{\mathfrak{D}}^{1}$
(4.1) $\theta=\sum_{I=1}^{{\mathrm{n}}}(a_{I}\operatorname{d}Q^{I}+b_{I}\operatorname{d}P^{I})+c\operatorname{d}\kappa,\ a_{I},~{}b_{I},~{}c\in\mathbb{R},~{}c\not=0,$
$\Omega$ is a 2-form with
(4.2) $\text{rank}(\Omega)=2{\mathrm{n}},$
and
(4.3) $\theta\wedge\Omega^{{\mathrm{n}}}\neq 0.$
We recall, see e.g. [73, §7] that if $A=a_{ij}\in M(n,\mathbb{C})$ then the vectorisation operator is defined as
${\mathrm{vec}}(A)^{t}=[a_{11},a_{21},\dots,a_{1n},a_{21},\dots,a_{2n},a_{31},\dots,a_{nn}]\in M(1,n^{2},\mathbb{C}),$
while the half-vectorisation operator is
${\mathrm{vech}}(A)^{t}=[a_{11},a_{12},\dots,a_{1n},a_{22},\dots,a_{2n},a_{33},\dots,a_{nn}]\in M(1,N_{1}),~{}N_{1}=\frac{n(n+1)}{2}.$
Note also that ${\mathrm{vech}}(A)=L_{n}{\mathrm{vec}}(A)$, where $L_{n}\in M(N_{1},n^{2})$ is the elimination matrix.
We endow the $n(n\!+\!3)\!+\!1$-dimensional manifold $\tilde{{\mathcal{X}}}^{J}_{n}$ with an ACOS structure $(\tilde{{\mathcal{X}}}^{J}_{n},\theta,\omega)$. For $\theta$ we take the formula (4.1), and consider $\Omega$ written in Darboux coordinates
(4.4) $\Omega=\sum_{I=1}^{{\mathrm{n}}}\operatorname{d}Q^{I}\wedge\operatorname{d}P^{I}.$
Lemma 2.
If in formula (3) of the Kähler-two form on ${\mathcal{D}}^{J}_{n}$ we make the partial the Cayley transform $(v,u)\rightarrow(v,\eta)=(x+\operatorname{i}y,q+\operatorname{i}p)$ (3), where $v\in MS(n,\mathbb{C})$, we get the Kähler  two-form $\omega_{{\mathcal{X}}^{J}_{n}}$
(4.5a) $\omega_{{\mathcal{X}}^{J}_{n}}(x,y,p,q)$ $=\omega_{1}+\omega_{2},$
(4.5b) $\omega_{1}$ $=\frac{k}{4}\operatorname{tr}(y^{-1}\operatorname{d}x\wedge y^{-1}\operatorname{d}y)=-\frac{k}{4}\operatorname{tr}(\operatorname{d}x\wedge\operatorname{d}y^{-1}).$
We arrange the elements of the symmetric matrices $x,y^{-1}$ in $\rm{vech}(x),\rm{vech}(y^{-1})$ s.t. there are $i$ elements on every row, $i=1$ on the last row, and $i=n$ on the first row. We use the notation
(4.6) $i_{d}:=N_{1}-\frac{i(i+1)}{2},\quad i=1,\dots,n.$
We vectorise $\omega_{1}$ (4.5b) as
(4.7) $\omega_{1}=\sum_{I=1}^{N_{1}}\operatorname{d}Q^{I}\wedge\operatorname{d}P^{I},$
where
(4.8) $(Q^{I},P^{I})=\left\{\begin{array}[]{ccc}(\frac{k}{4}x_{i_{d}+1,i_{d}+1},&-{y}^{-1}_{i_{d}+1,i_{d}+1})&,I=i_{d}+1;\\ (\frac{k}{2}x_{i_{d}+1,i_{d}+j},&-{y}^{-1}_{i_{d}+1,i_{d}+j})&,j=2,\dots,i,I=i_{d}+2,\dots,i_{d}+i.\end{array}\right.$
For $\omega_{2}$ in (4.5a) we have
(4.9) $\omega_{2}=2\nu\operatorname{d}Q\wedge\operatorname{d}P^{t}=\sum_{I=N_{1}+1}^{N_{1}+n}\operatorname{d}Q^{i}\wedge\operatorname{d}P^{i},~{}Q^{N_{1}+i}=2\nu q_{i},P^{N_{1}+i}=p_{i},~{}i=1,\dots,n.$
Proof.
With
$H=-\frac{1}{2\operatorname{i}}y^{-1}\operatorname{d}(x+\operatorname{i}y).$
we get
$\operatorname{tr}[H\wedge\bar{H}]=\frac{1}{4}y^{-1}_{ik}(\operatorname{d}x_{kj}+\operatorname{i}\operatorname{d}y_{kj})\wedge y^{-1}_{jl}(\operatorname{d}x_{li}-\operatorname{i}\operatorname{d}y_{li})=\frac{\operatorname{i}}{2}\operatorname{tr}(y^{-1}\operatorname{d}y\wedge y^{-1}\operatorname{d}x).$
Now we calculate $\omega_{2}$. Using the symmetry of the matrices $x,y$, we get successively
$\operatorname{tr}(G^{t}D\wedge\bar{G})$ $=\frac{\operatorname{i}}{2}\operatorname{tr}\{[\operatorname{d}p(x+\operatorname{i}y)+\operatorname{d}q]y^{-1}\wedge[(x-\operatorname{i}y)\operatorname{d}p^{t}+\operatorname{d}q^{t}]$
$=\frac{\operatorname{i}}{2}\operatorname{tr}\{\operatorname{d}p(x\!+\!\operatorname{i}y)y^{-1}\wedge[\!(x\!-\!\operatorname{i}y)\operatorname{d}p^{t}\!+\!\operatorname{d}q^{t}\!]\!+\!\operatorname{d}qy^{-1}\wedge[(\!x-\!\operatorname{i}y\!)\operatorname{d}p^{t}\!+\!\operatorname{d}q^{t}\!]$
$=\frac{\operatorname{i}}{2}\operatorname{tr}\{\operatorname{d}p[(xy^{-1}x+yy^{-1}y)\wedge\operatorname{d}p^{t}+(xy^{-1}+\operatorname{i})\wedge\operatorname{d}q^{t}]$
$+\operatorname{d}q[(y^{-1}x-\operatorname{i})\wedge\operatorname{d}p^{t}+q\wedge y^{-1}\operatorname{d}q^{t}]\}$
$=\frac{\operatorname{i}}{2}\operatorname{tr}\{\operatorname{d}p(xy^{-1}x+yy^{-1}y)\operatorname{d}p^{t}+\operatorname{d}pxy^{-1}\wedge\operatorname{d}q^{t}$
$+\operatorname{d}qy^{-1}\wedge\operatorname{d}q^{t}+\operatorname{i}(\operatorname{d}p\wedge\operatorname{d}q^{t}-\operatorname{d}q\wedge\operatorname{d}p^{t})\}$
$=\frac{\operatorname{i}}{2}\operatorname{tr}\{\operatorname{d}p(xy^{-1}x\!+\!yy^{-1}y)\wedge\operatorname{d}p^{t}+\operatorname{d}qy^{-1}\wedge\operatorname{d}q^{t}$
$+\operatorname{d}pxy^{-1}\wedge\operatorname{d}q^{t}+\operatorname{d}qy^{-1}x\wedge\operatorname{d}p^{t}\!+\!2\operatorname{i}\operatorname{d}p\wedge\operatorname{d}q^{t}\}$
$=-\operatorname{d}p\wedge\operatorname{d}q^{t},$
and we find
(4.11) $\omega_{2}=2\nu\operatorname{d}q^{t}\wedge\operatorname{d}p.$
We identify $\Omega$ in formula (4.4) with $\omega$ from (4.5a), where $\omega_{1}$ $(\omega_{2})$ is given by (4.7), (4.8) (respective(4.9)). We find:
(4.12a) $a_{I}$ $=b_{I}=0,I=1,\dots,N_{1};$
(4.12b) $a_{N_{1}+i}$ $=-\frac{\sqrt{\delta}}{2\nu}p_{i};~{}~{}b_{N_{1}+i}=\sqrt{\delta}q_{i},i=1,\dots,n;$
(4.12c) $c$ $=\sqrt{\delta},~{}~{}~{}{\mathrm{n}}=\frac{n(n+3)}{2}.$
Now we identify $\lambda_{6}$ in formula (2.51) with $\theta$ given by (4.1). We write $\lambda_{6}$ as
$\lambda_{6}=\sqrt{\delta}\{[\sum_{I=1}^{N_{1}}-P_{I}\operatorname{d}Q_{I}+Q_{I}\operatorname{d}P_{I}]+\sum_{i=1}^{n}[-P_{N_{1}+i}\operatorname{d}Q_{N_{1}+i}+Q_{N_{1}+i}\operatorname{d}P_{N_{1}+i}]+\operatorname{d}\kappa\},$
and $\theta$ as
$\theta=\sum_{I=1}^{N_{1}}[a_{I}\operatorname{d}Q_{I}+b_{I}\operatorname{d}P_{I}]+\sum_{I=N_{1}+1}^{{\mathrm{n}}}[a_{I}\operatorname{d}Q_{I}+b_{I}\operatorname{d}P_{I}]+c\operatorname{d}\kappa.$
[…]
We have
$\operatorname{d}\omega=0,\quad\theta\wedge\omega^{\rm{n}}=c\frac{{\mathrm{n}}}{2}\Pi_{I=1}^{\frac{{\mathrm{n}}}{2}}\operatorname{d}Q^{I}\wedge\operatorname{d}P^{I}\wedge\operatorname{d}\kappa,$
i.e. condition (4.3) is satisfied because of (4.2). ∎
We endow $\tilde{{\mathcal{X}}}^{J}_{1}$ with a generalized transitive almost cosymplectic (GTACOS) structure [30], i.e. an ACOS structure $(M,\theta,\Omega)$ such that
$\operatorname{d}\Omega=0.$
Lemma 3, extracted from [30, Lemma 1], is a particular case of Lemma 2. To proof (2.25b) we introduce the relations (2.24a) into (1).
Alternatively, we introduce in formula (2.11) of ${\mathcal{A}}(w,z)$ $z=\eta-w\bar{\eta}$ (2.12) and we find
(4.13) ${\mathcal{A}}\wedge\bar{{\mathcal{A}}}=P\operatorname{d}\eta\wedge\operatorname{d}\bar{\eta}.$
Introducing (2.27), (3.29) and (4.13), we get again (2.25b).
Lemma 3.
If we introduce into the Kähler two-form $\omega_{{\mathcal{D}}^{J}_{1}}(w,z)$ (2.11) the second partial Cayley transform (2.24a) $(w,z)\rightarrow(x,y,q,p),y>0$ we get the symplectic two-form (2.25b).
In the notation of [27], we introduce on the extended 5-dimensional Siegel-Jacobi half-plane $\tilde{{\mathfrak{X}}}^{J}_{1}$ parametrized in $(x,y,p,q,\kappa)$ the almost cosymplectic structure $(\tilde{{\mathcal{X}}}^{J}_{1},\theta,\omega)$, where $\theta=\lambda_{6}$ and $\omega$ is (2.25b), i.e.
(4.14a) $\theta$ $=\sqrt{\delta}(\operatorname{d}\kappa-p\operatorname{d}q+q\operatorname{d}p),\quad\delta>0,$
(4.14b) $\omega$ $=\frac{k}{y^{2}}\operatorname{d}x\wedge\operatorname{d}y+2\nu\operatorname{d}q\wedge\operatorname{d}p,\quad y>0.$
We have
(4.15) $\operatorname{d}\omega=0,\quad\theta\wedge\omega^{2}=4\frac{k\nu\sqrt{\delta}}{y^{2}}\operatorname{d}x\wedge\operatorname{d}y\wedge\operatorname{d}q\wedge\operatorname{d}p\wedge\operatorname{d}\kappa,$
and $(\tilde{{\mathfrak{X}}}^{J}_{1},\theta,\omega)$ verifies the condition [30, (5.5)] of an almost cosymplectic manifold.
With a formula of the type (3.2b), the Kähler two-form (4.14b) on ${\mathcal{X}}^{J}_{1}$ corresponds to the Kähler  potential
(4.16) $f(v,\bar{v},\eta,\bar{\eta})=-2k\log\frac{v-\bar{v}}{2\operatorname{i}}+\nu[\eta\bar{\eta}+g(\eta)+h(\bar{\eta})],$
in particular we get
(4.17) $f(v,\bar{v},\eta,\bar{\eta})=-2k\log\frac{v-\bar{v}}{2\operatorname{i}}-\frac{\nu}{2}(\eta-\bar{\eta})^{2},$
which correspond to particular values of functions in (3.7a).
If in (4.17) we make the change of coordinates (2.24b) in (4.17), we get
(4.18) $f(w,\bar{w},\eta,\bar{\eta})=-2k\log\frac{1-w\bar{w}}{(1-w)(1-\bar{w})}-\frac{\nu}{2}(\eta-\bar{\eta})^{2},$
which corespond to the particular values $f(w)=2k\log\frac{1-w}{w}$, $g(\bar{w})=2k\log(1-\bar{w})$, $f^{\prime}(\eta)={g^{\prime}(\bar{\eta})}=-\frac{\nu}{2}\eta^{2}$ in (3.7b).
If we apply to the reproducing Kernel (4.18) a formula of the type (3.1), we get again (2.14a).
In Darboux coordinates we have a particular almost cosymplectic manifold $(\tilde{{\mathfrak{X}}}^{J}_{1},\theta,\omega)$ verifying [30, (5.5)] and in addition the condition
$\operatorname{d}\omega=0.$
$(\tilde{{\mathfrak{X}}}^{J}_{1},\theta,\omega)$ was called generalized transitive almost cosymplectic manifold [30, GTACOS].
4.2. Connection matrix on $\tilde{{\mathcal{X}}}^{J}_{1}$
We determined the Christofell’s symbols corresponding to the Riemannian metric (2.51) of the extended Siegel-Jacobi upper half-plane $\tilde{{\mathcal{X}}}^{J}_{1}$ [29, page 22]. In formulas below we have included only the $\Gamma$-s which are not given in (3.45)
(4.19) $\begin{array}[]{llll}\!\Gamma^{p}_{pp}\!=\!2\tau\frac{xq}{y}&\!\Gamma^{p}_{pq}\!=\!\tau\frac{q\!-\!px}{y}&\!\Gamma^{p}_{p\kappa}\!=\!\tau\frac{x}{y}&\!\Gamma^{p}_{qq}\!\!=\!\!-\!2\tau\frac{p}{y}\!~{}\!\Gamma^{p}_{q\kappa}\!=\!\tau\frac{1}{y}\\ \!\Gamma^{q}_{pp}\!=\!-2\tau\frac{qS}{y}&\!\Gamma^{q}_{pq}\!\!=\!\!\tau\frac{-xq\!\!+\!\!pS}{y}&\!\Gamma^{q}_{p\kappa}\!=\!-\!\tau\frac{S}{y}&\!\Gamma^{q}_{qq}\!=\!2\tau\!\frac{xp}{y}\!~{}\!\Gamma^{q}_{q\kappa}\!=\!-\!\tau\frac{x}{y}\\ \!\Gamma^{\kappa}_{xp}\!=\!\frac{py^{2}\!-\!x\xi}{2y^{2}}&\!\Gamma^{\kappa}_{xq}\!=\!-\!\frac{\xi}{2y^{2}}&\Gamma^{\kappa}_{yp}\!\!=\!\!-\!\frac{2px+q}{2y}&\!\Gamma^{\kappa}_{yq}\!=\!-\frac{p}{2y}~{}\Gamma^{\kappa}_{pp}\!\!=\!\!-\!2\!\tau\!\frac{q}{y}(\!p\!S\!+\!q\!x\!)\\ \!\Gamma^{\kappa}_{pq}\!=\!\tau\frac{p^{2}S-q^{2}}{y}&\!\Gamma^{\kappa}_{p\kappa}\!\!=\!-\!\tau\frac{pS+qx}{y}&\Gamma^{\kappa}_{qq}\!\!=\!\!2\tau\frac{p\xi}{y}&\!\Gamma^{\kappa}_{q\kappa}\!\!=-\!\tau\frac{\xi x}{y}.\end{array}$
where
$\tau:=\frac{\delta}{\gamma},\quad\xi:=px+q.$
We determine the connection matrix on the extended Siegel-Jacobi upper half-plane in the S-coordinates $(x,y,q,p,\kappa)$
(4.20k) $\!\theta^{\prime}_{\tilde{{\mathcal{X}}}^{J}_{1}}(x,y,q,p,\kappa)\!=$ $\left(\begin{array}[]{ccccc}\theta^{\prime x}_{x}&\theta^{\prime x}_{y}&\theta^{\prime x}_{q}&\theta^{\prime x}_{p}&\theta^{\prime x}_{\kappa}\\ \theta^{\prime y}_{x}&\theta^{\prime y}_{y}&\theta^{\prime y}_{q}&\theta^{\prime y}_{p}&\theta^{\prime y}_{\kappa}\\ \theta^{\prime q}_{x}&\theta^{\prime q}_{y}&\theta^{\prime q}_{q}&\theta^{\prime q}_{p}&\theta^{\prime q}_{\kappa}\\ \theta^{\prime p}_{x}&\theta^{\prime p}_{y}&\theta^{\prime p}_{q}&\theta^{\prime p}_{p}&\theta^{\prime p}_{\kappa}\\ \theta^{\prime\kappa}_{x}&\theta^{\prime\kappa}_{y}&\theta^{\prime\kappa}_{q}&\theta^{\prime\kappa}_{p}&\theta^{\prime\kappa}_{\kappa}\end{array}\right)~{}~{}$ $=\left(\begin{array}[]{ccccc}\theta^{x}_{x}&\theta^{x}_{y}&\theta^{x}_{q}&\theta^{x}_{p}&0\\ \theta^{y}_{x}&\theta^{y}_{y}&\theta^{y}_{q}&\theta^{y}_{p}&0\\ \theta^{q}_{x}&\theta^{q}_{y}&\theta^{\prime q}_{q}&\theta^{\prime q}_{p}&\theta^{\prime q}_{\kappa}\\ \theta^{p}_{x}&\theta^{p}_{y}&\theta^{\prime p}_{q}&\theta^{\prime p}_{p}&\theta^{\prime p}_{\kappa}\\ \theta^{\prime\kappa}_{x}&\theta^{\prime\kappa}_{y}&\theta^{\prime\kappa}_{q}&\theta^{\prime\kappa}_{p}&\theta^{\prime\kappa}_{\kappa}\end{array}\right).$
With (3.45) and (4.19), we find for the matrix elements of (4.2) the values
(4.21r) $\begin{array}[]{ll}\theta^{\prime q}_{q}&=\Gamma^{q}_{qx}\operatorname{d}x+\Gamma^{q}_{qy}\operatorname{d}y+\Gamma^{q}_{qq}\operatorname{d}q+\Gamma^{q}_{qp}\operatorname{d}p+\Gamma^{q}_{q\kappa}\operatorname{d}\kappa\\ &=-\frac{x}{2y^{2}}\operatorname{d}x-\frac{1}{2y}\operatorname{d}y+\frac{\tau}{y}\left(2xp\operatorname{d}q+(-xq+pS)\operatorname{d}p-x\operatorname{d}\kappa\right)\\ \theta^{\prime q}_{p}&=\theta^{q}_{p}+\Gamma^{q}_{qp}\operatorname{d}p+\Gamma^{q}_{q\kappa}\operatorname{d}\kappa\\ &=\frac{y^{2}-x^{2}}{2y^{2}}\operatorname{d}x-\frac{x}{y}\operatorname{d}y+\frac{\tau}{y}\left((-xq+pS)\operatorname{d}q-2qS\operatorname{d}p-x\operatorname{d}\kappa)\right)\\ \theta^{\prime q}_{\kappa}&=\Gamma^{\prime q}_{\kappa q}\operatorname{d}q+\Gamma^{q}_{\kappa p}\operatorname{d}p=-\tau\frac{x}{y}\operatorname{d}q-\tau\frac{x}{y}\operatorname{d}p\\ \theta^{\prime p}_{q}&=\Gamma^{p}_{qx}\operatorname{d}x+\Gamma^{p}_{qq}\operatorname{d}q+\Gamma^{p}_{qp}\operatorname{d}p+\Gamma^{p}_{q\kappa}\operatorname{d}\kappa\\ &=\frac{x}{2y^{2}}\operatorname{d}x+\frac{\operatorname{d}y}{2y}+\frac{\tau}{y}\left(-2p\operatorname{d}q+(q-px)\operatorname{d}p+\operatorname{d}\kappa\right)\\ \theta^{\prime p}_{p}&=\Gamma^{p}_{px}\operatorname{d}x\!+\!\Gamma^{p}_{py}\operatorname{d}y\!+\!\Gamma^{p}_{pq}\operatorname{d}q\!+\!\Gamma^{p}_{pp}\operatorname{d}p\!+\!\Gamma^{p}_{p\kappa}\operatorname{d}\kappa\\ &=\frac{x}{2y^{2}}\operatorname{d}x\!+\!\frac{x}{2y}\operatorname{d}y\!+\!\frac{\tau}{y}\left((q-px)\operatorname{d}q+2xq\operatorname{d}p+x\operatorname{d}\kappa\right)\\ \theta^{\prime p}_{\kappa}&=\Gamma^{p}_{\kappa q}\operatorname{d}q+\Gamma^{p}_{\kappa p}\operatorname{d}p=\frac{\tau}{y}(\operatorname{d}q+x\operatorname{d}p)\\ \theta^{\prime\kappa}_{x}&=\Gamma^{\kappa}_{xq}\operatorname{d}q+\Gamma^{\kappa}_{x\kappa}\operatorname{d}p=-\frac{\xi}{2y^{2}}\operatorname{d}q+\frac{py^{2}-x\xi}{2y^{2}}\operatorname{d}p\\ \theta^{\prime\kappa}_{y}&=\Gamma^{\kappa}_{yq}\operatorname{d}q+\Gamma^{\kappa}_{yp}\operatorname{d}p=-\frac{p}{2y}\operatorname{d}q-\frac{2px+q}{2y}\operatorname{d}p\\ \theta^{\prime\kappa}_{q}&=\Gamma^{\kappa}_{qx}\operatorname{d}x+\Gamma^{\kappa}_{qy}\operatorname{d}y+\Gamma^{\kappa}_{qq}\operatorname{d}q+\Gamma^{\kappa}_{qp}\operatorname{d}p+\Gamma^{\kappa}_{q\kappa}\operatorname{d}\kappa\\ &=-\frac{\xi}{2y^{2}}\operatorname{d}x-\frac{p}{2y}\operatorname{d}y+\frac{\tau}{y}(2p\xi\operatorname{d}q+(p^{2}S-q^{2})\operatorname{d}p-\xi x\operatorname{d}\kappa)\\ \theta^{\prime\kappa}_{p}&=\Gamma^{\kappa}_{px}\operatorname{d}x+\Gamma^{\kappa}_{py}\operatorname{d}y+\Gamma^{\kappa}_{pq}\operatorname{d}q+\Gamma^{\kappa}_{pp}\operatorname{d}p+\Gamma^{\kappa}_{p\kappa}\operatorname{d}\kappa\\ &=\frac{py^{2}-x\xi}{2y^{2}}\operatorname{d}x-\frac{2px+q}{2y}\operatorname{d}y+\frac{\tau}{y}((p^{2}S-q^{2})\operatorname{d}q-2q(pS+qx)\operatorname{d}p-(pS+qx)\operatorname{d}\kappa)\\ \theta^{\prime\kappa}_{\kappa}&=\Gamma^{\kappa}_{\kappa q}\operatorname{d}q+\Gamma^{\kappa}_{\kappa p}\operatorname{d}p=-\frac{\tau}{y}(\xi x\operatorname{d}q+(pS+qx)\operatorname{d}p)\end{array}$
4.3. Covariant derivative of one-forms on ${\mathcal{X}}^{J}_{1}$ and $\tilde{{\mathcal{X}}}^{J}_{1}$
The covariant derivative of a contravariant vector (one-form) is given by
(4.22) $Du_{i}=-\theta^{i}_{j}u_{j}=-u_{j}\Gamma^{i}_{jk}u_{k}=-u_{j}\Gamma^{i}_{kj}u_{k}.$
The covariant derivative of $\operatorname{d}z$ on ${\mathcal{D}}^{J}_{1}$ has the expression [25, (41)]
(4.23) $D(\operatorname{d}z)=\left(\begin{array}[]{cc}\operatorname{d}z\operatorname{d}w\end{array}\right)\left(\begin{array}[]{cc}\lambda\bar{\eta}&\lambda\bar{\eta}^{2}-\frac{\bar{w}}{P}\\ \lambda\bar{\eta}^{2}-\frac{\bar{w}}{P}&\lambda\bar{\eta}^{3}\end{array}\right)\left(\begin{array}[]{c}\operatorname{d}z\\ \operatorname{d}w\end{array}\right).$
The covariant derivative of $\operatorname{d}w$ has the expression [25, (42)]
(4.24) $-D(\operatorname{d}w)=\left(\begin{array}[]{cc}\operatorname{d}z\operatorname{d}w\end{array}\right)\left(\begin{array}[]{cc}\lambda&\lambda\bar{\eta}\\ \lambda\bar{\eta}&\lambda\bar{\eta}^{2}+2\frac{\bar{w}}{P}\end{array}\right)\left(\begin{array}[]{c}\operatorname{d}z\\ \operatorname{d}w\end{array}\right).$
We calculate the covariant derivative of the S-variables $x,y,p,q$ on ${\mathcal{X}}^{J}_{1}$
(4.25) $Dx=\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\end{array}\right)^{t}\left(\begin{array}[]{cccc}0&\frac{1}{y}&0&0\\ \frac{1}{y}&0&0&0\\ 0&0&0&\frac{\epsilon\,}{2}y\\ 0&0&\frac{\epsilon\,}{2}y&\epsilon\,xy\end{array}\right)\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\end{array}\right).$
(4.26) $Dy=\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\\ \end{array}\right)^{t}\left(\begin{array}[]{cccc}-\frac{1}{y}&0&0&0\\ 0&\frac{1}{y}&0&0\\ 0&0&-\frac{\epsilon\,}{2}&-\frac{\epsilon\,}{2}x\\ 0&0&-\frac{\epsilon\,}{2}x&\frac{\epsilon\,}{2}(y^{2}-x^{2})\end{array}\right)\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\end{array}\right).$
(4.27) $Dq=\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\end{array}\right)^{t}\left(\begin{array}[]{cccc}0&0&\frac{x}{2y^{2}}&\frac{x^{2}-y^{2}}{2y^{2}}\\ 0&\frac{x}{y}&0&0\\ \frac{x}{2y^{2}}&0&0&0\\ \frac{x^{2}-y^{2}}{2y^{2}}&0&0&0\end{array}\right)\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\end{array}\right).$
(4.28) $Dp=\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\\ \end{array}\right)^{t}\left(\begin{array}[]{cccc}0&0&-\frac{1}{2y^{2}}&-\frac{x}{2y^{2}}\\ 0&-\frac{1}{2y}&0&0\\ -\frac{1}{2y^{2}}&0&0&0\\ -\frac{x}{2y^{2}}&0&0&0\end{array}\right)\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\end{array}\right).$
We calculate the covariant derivative of the S-variables $x,y,p,q,\kappa$ on $\tilde{{\mathcal{X}}}^{J}_{1}$
(4.29) $Dx=\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\\ \operatorname{d}\kappa\end{array}\right)^{t}\left(\begin{array}[]{ccccc}0&\frac{1}{y}&0&0&0\\ \frac{1}{y}&0&0&0&0\\ 0&0&0&\frac{\epsilon\,}{2}y&0\\ 0&0&\frac{\epsilon\,}{2}{y}&\epsilon\,xy&0\\ 0&0&0&0&0\end{array}\right)\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\\ \operatorname{d}\kappa\end{array}\right).$
(4.30) $Dy=\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\\ \operatorname{d}\kappa\end{array}\right)^{t}\left(\begin{array}[]{ccccc}-\frac{1}{y}&0&0&0&0\\ 0&\frac{1}{y}&0&0&0\\ 0&0&-\frac{\epsilon\,}{2}&-\frac{\epsilon\,}{2}x&0\\ 0&0&-\frac{\epsilon\,}{2}x&\frac{\epsilon\,}{2}(y^{2}-x^{2})&0\\ 0&0&0&0&0\end{array}\right)\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\\ \operatorname{d}\kappa\end{array}\right).$
(4.31) $Dq=\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\\ \operatorname{d}\kappa\end{array}\right)^{t}\left(\begin{array}[]{ccccc}0&0&\frac{x}{2y^{2}}&\frac{x^{2}-y^{2}}{2y^{2}}&0\\ 0&0&\frac{1}{2y}&\frac{x}{y}&0\\ \frac{x}{2y^{2}}&\frac{1}{2y}&-2\frac{\tau}{y}xp&\frac{\tau}{y}(xq-pS)&\frac{\tau}{y}x\\ \frac{x^{2}-y^{2}}{2y^{2}}&\frac{x}{y}&\frac{\tau}{y}(xq-pS)&\frac{\tau}{y}2qS&\frac{\tau}{y}S\\ 0&0&\frac{\tau}{y}x&\frac{\tau}{y}S&0\end{array}\right)\left(\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\\ \operatorname{d}\kappa\end{array}\right).$
(4.32) $Dp\!=\!\left(\!\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\!\operatorname{d}\kappa\end{array}\right)^{t}\left(\!\begin{array}[]{ccccc}0&0&-\frac{1}{2y^{2}}&-\frac{x}{2y^{2}}&0\\ 0&0&0&-\frac{1}{2y}&0\\ -\frac{1}{2y^{2}}&0&2\frac{\tau}{y}p&\frac{\tau}{y}(q-px)&-\frac{\tau}{y}\\ -\frac{x}{2y^{2}}&-\frac{1}{2y}&\frac{\tau}{y}(q-px)&-\tau\frac{xq}{y}&-\frac{\tau}{y}x\\ 0&0&-\frac{\tau}{y}&-\frac{\tau}{y}x&0\end{array}\right)\!\left(\!\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\\ \operatorname{d}\kappa\end{array}\right)\!$
(4.33) $D\kappa\!=\!\left(\!\begin{array}[]{c}\operatorname{d}x\\ \operatorname{d}y\\ \operatorname{d}q\\ \operatorname{d}p\\ \operatorname{d}\kappa\end{array}\right)^{t}\!\left(\!\begin{array}[]{ccccc}\!0&\!0&\!\frac{\xi}{2y^{2}}&\!\frac{x\xi-py^{2}}{2y^{2}}&\!\!0\\ \!0&\!\!0&\!\frac{p}{2y}&\!\frac{2px+q}{2y}&\!\!0\\ \!\frac{\xi}{2y^{2}}&\!\!\frac{p}{2y}&\!-2\frac{\tau}{y}p\xi&\!-\frac{\tau}{y}(p^{2}S\!-\!q^{2})&\!\!\frac{\tau}{y}\xi x\\ \!\frac{x\xi-py^{2}}{2y^{2}}&\!\!\frac{2px+q}{2y}&\!\frac{\tau}{y}(q^{2}-p^{2}S)&2\frac{\tau q}{y}(pS+qx)&\!\!\frac{\tau}{y}(pS\!+\!qx)\\ \!0&\!\!0&\!\frac{\tau}{y}\xi x&\frac{\tau}{y}(pS\!+\!qx)&\!\!0\end{array}\right)\!\left(\!\begin{array}[]{c}\!\operatorname{d}x\\ \operatorname{d}y\\ \!\operatorname{d}q\\ \!\operatorname{d}p\\ \!\operatorname{d}\kappa\end{array}\right)\!$
5. Appendix
5.1. Theory
5.1.1. Connections on real manifolds
a) Connections on vector bundles. Following [49, page 101], let $E$ be a $q$-dimensional real vector bundle with projection $\psi:E\rightarrow M$ on the $m$-dimensional manifold $M$ and let $\Gamma(E)$ be the set of smooth sections of $E$ on $M$.
Definition 1.
A connection on the vector bundle $E$ is a map
$D:\Gamma(E)\rightarrow\Gamma(T^{*}(M)\otimes E)$
such that
(5.1a) $1$ $~{}~{}D(s_{1}+s_{2})=Ds_{1}+Ds_{2},\quad\forall s_{1},s_{2}\in\Gamma(E),$
(5.1b) $2$ $~{}~{}D(\alpha s)=\operatorname{d}\alpha\otimes s+\alpha Ds,\quad\alpha\in C^{\infty}(M).$
The absolute differential quotient or the covariant derivative of the section $s$ along $X\in{\mathfrak{D}}^{1}$ is defined as
(5.2) $D_{X}s:=<X,Ds>,$
where <,><,>< , > is the pairing between the tangent space $T(M)$ and the cotangent space $T^{*}(M)$.
Chose a local field frame of $E$ on the neighborhood $U\subset M$, i.e. $q$-linearly independent smooth sections $s_{\alpha}$. Then at every point $p\in U$$\operatorname{d}u^{i}\otimes s_{\alpha},1\leq i\leq m;1\leq\alpha\leq q$} form a basis of $T^{*}_{p}\otimes E$ and
(5.3) $Ds_{\alpha}=\sum_{\beta=1}^{q}\omega^{\beta}_{\alpha}\otimes s_{\beta},\quad\omega_{\alpha}^{\beta}=\sum_{1\leq i\leq m}\Gamma^{\beta}_{\alpha i}\operatorname{d}u^{i},$
where $\omega^{\beta}_{\alpha}$ are real valued 1-forms on $U$ and $\Gamma^{\beta}_{\alpha i}$ are smooth functions on U. Sometimes instead of $\omega$ it is used the symbol $\theta$, see (3.11), (5.37), (5.48).
If we denote
(5.4) $S=\left(\begin{array}[]{c}s_{1}\\ \vdots\\ s_{q}\end{array}\right),\quad\omega=\left(\begin{array}[]{ccc}\omega^{1}_{1}&\cdots&\omega^{q}_{1}\\ \vdots&\ddots&\vdots\\ \omega^{1}_{q}&\cdots&\omega^{q}_{q}\end{array}\right),$
then (5.3) can be written as
(5.5) $DS=\omega\otimes S.$
The matrix of real one-forms $\omega$ in (5.4) is called the connection matrix or connection form [53].
If $S^{\prime}=(s^{\prime}_{1},\dots,s^{\prime}_{q})^{t}$ is another local frame field on $U$ and
(5.6) $S^{\prime}=AS,\quad A=\left(\begin{array}[]{ccc}a^{1}_{1}&\cdots&a^{q}_{1}\\ \vdots&\ddots&\vdots\\ a^{1}_{q}&\cdots&a^{q}_{q}\end{array}\right),\quad A\in M(q,\mathbb{R}),$
then
(5.7) $DS^{\prime}=\omega^{\prime}\otimes S^{\prime}.$
With (5.5), (5.6), we got
$\operatorname{d}A\otimes S+ADS=\omega^{\prime}\otimes AS,$
and finally
(5.8) $\omega^{\prime}=\operatorname{d}A\cdot A^{-1}+A\cdot\omega\cdot A^{-1}.$
Definition 2.
The $q\times q$ matrix of two-forms
(5.9) $\Omega=\operatorname{d}\omega-\omega\wedge\omega$
is called the curvature matrix of the connection $D$ on $U$ [49, Definition 1.2 page 108]. Sometimes the curvature matrix $\Omega$ is denoted $\Theta$, see (5.37) below.
In the new system of coordinate $S^{\prime}$ (5.6) the curvature matrix $\Omega^{\prime}$ is
(5.10) $\Omega^{\prime}=A\cdot\Omega\cdot A^{-1}.$
The curvature matrix $\Omega$ defines a linear transformation from $\Gamma(E)$ to $\Gamma(E)$. For
(5.11) $\Gamma(E)\ni s=\sum_{\alpha=1}^{q}\lambda^{\alpha}s_{\alpha}.$
let $X,Y\in{\mathfrak{D}}^{1}$, and define
(5.12) $R(X,Y)s:=\sum_{\alpha=1}^{q}\lambda^{\alpha}<X\wedge Y,\Omega^{\beta}_{\alpha}>s_{\beta}|_{p}.$
$R(X,Y)$ is called the curvature operator of the connection $D$ and [49, Theorem 1.3 page 109]
$R(X,Y)=D_{X}D_{Y}-D_{Y}D_{X}-D_{[X,Y]}.$
We have [49, page 117]
(5.13) $\Omega^{j}_{i}=\frac{1}{2}R^{j}_{ikl}\operatorname{d}u^{k}\wedge\operatorname{d}u^{l},\quad R^{j}_{ikl}=\frac{\partial\Gamma^{j}_{il}}{\partial u^{k}}-\frac{\partial\Gamma^{j}_{ik}}{\partial u^{l}}+\Gamma^{h}_{il}\Gamma^{j}_{hk}-\Gamma^{h}_{ik}\Gamma^{j}_{hl},$
and the $(1,3)$-tensor
$R=R^{j}_{ikl}\frac{\partial}{\partial u^{j}}\otimes\operatorname{d}u^{i}\otimes\operatorname{d}u^{k}\otimes\operatorname{d}u^{l}$
is called the curvature tensor of the affine connection $D$.
The section $s$ of a vector bundle $E$ is called parallel section if [49, page 116]
(5.14) $Ds=0.$
For (5.11), equation (5.14) with (5.2), (1) becomes
(5.15) $\operatorname{d}\lambda^{\alpha}+\sum_{\beta=1}^{q}\lambda^{\beta}\omega^{\alpha}_{\beta}=0,\quad 1\leq\alpha\leq q.$
Definition 3.
Let $C$ be a parametrized curve and $X$ a tangent vector field along $C$. If the section $s$ of the vector bundle $E$ on $C$ satisfies
(5.16) $D_{X}s=0,$
then $s$ is said parallel along the curve $C$.
If
(5.17) $X(t)=\sum_{i=1}^{m}\frac{\operatorname{d}u^{i}}{\operatorname{d}t}\left(\frac{\partial}{\partial u^{i}}\right)_{C(t)},$
and $s$ is given by (5.11) then (5.16) reads
(5.18) $\frac{\operatorname{d}\lambda^{\alpha}}{\operatorname{d}t}+\sum_{\beta,i}\Gamma^{\alpha}_{\beta i}\frac{\operatorname{d}u^{i}}{\operatorname{d}t}\lambda^{\beta}=0,\quad 1\leq\alpha\leq q.$
If any vector $v\in E_{p}$ is given at a point $p$ on $C$, then it determines uniquely a vector field along $C$, called parallel displacement of $v$ along $C$.
b) Affine connections. A connection on the $m$-dimensional tangent vector bundle $T(M)$ is called affine connection on $M$ [49, §  4-2].
Formulas (5.3) in the natural basis {$s_{i}=\frac{\partial}{\partial u^{i}},1\leq i\leq m$} in the local coordinate system $(U,u^{i})$ on $M$ became
(5.19) $Ds_{i}=\omega^{j}_{i}\otimes s_{j}=\Gamma^{j}_{ik}\operatorname{d}u^{k}\otimes s_{j},$
and the smooth functions $\Gamma^{j}_{ik}$ on $U$ are called coefficients of the connection $D$ with respect to the local coordinates $u^{i}$.
In [60, §4] the affine connection $D$ is denoted $\nabla$, $D_{X}$ defined at (5.2) is denoted $\nabla_{X}$ and (1) becomes
(5.20a) $1$ $\nabla_{fX+gY}=f\nabla_{X}+g\nabla_{Y},~{}X,Y\in{\mathfrak{D}}^{1},~{}f,g\in C^{\infty}(M),$
(5.20b) $2$ $\nabla_{X}(fY)=f\nabla_{X}(Y)+(Xf)Y,\quad Xf=<X,\operatorname{d}f>.$
Equation (5.2) for $X=s_{l}$, $s=s_{i}$ becomes [60, (1), page 27]
(5.21) $\nabla_{\frac{\partial}{\partial u^{l}}}\left(\frac{\partial}{\partial u^{i}}\right)=\Gamma^{j}_{il}\frac{\partial}{\partial u^{j}}.$
If $(W,w^{i})$ is another coordinate system of $M$ and $s^{\prime}_{i}=\frac{\partial}{\partial w^{i}}$, then on $U\cap W\not=0$ we have
(5.22) $S^{\prime}=J_{WU}\cdot S,\quad J_{WU}=\left(\begin{array}[]{ccc}\frac{\partial u^{1}}{\partial w^{1}}&\cdots&\frac{\partial u^{m}}{\partial w^{1}}\\ \vdots&\ddots&\vdots\\ \frac{\partial u^{1}}{\partial w^{m}}&\cdots&\frac{\partial u^{m}}{\partial w^{m}}\end{array}\right),$
and (5.7) becomes [49, pages 113, 114]
(5.23a) $\omega^{\prime}$ $=\operatorname{d}J_{WU}\cdot J^{-1}_{WU}+J_{WU}\cdot\omega\cdot J^{-1}_{WU},$
(5.23b) $\omega^{\prime j}_{i}$ $=\operatorname{d}\left(\frac{\partial u^{p}}{\partial w^{i}}\right)\frac{\partial w^{j}}{\partial u^{p}}+\frac{\partial u^{p}}{\partial w^{i}}\frac{\partial w^{j}}{\partial u^{q}}w^{q}_{p},$
(5.23c) $\Gamma^{\prime j}_{ik}$ $=\Sigma_{pqr}\Gamma^{q}_{pr}\frac{\partial w^{j}}{\partial u^{q}}\frac{\partial u^{p}}{\partial w^{i}}\frac{\partial u^{r}}{\partial w^{k}}+\Sigma_{p}\frac{\partial^{2}u^{p}}{\partial w^{i}\partial w^{k}}\cdot\frac{\partial w^{j}}{\partial u^{p}},$
where $\omega^{\prime j}_{i}=\Gamma^{\prime j}_{ik}\operatorname{d}\omega^{k}$. Fomula (5.23c) appears also as [60, (2) page 27]. $DX$ is a (1,1)-type tensor field on $M$, called the absolute differential of $X$, [49, page 114].
Let $T^{r}_{s}$ be the tensor product of the tangent and cotangent bundles. If $t$ is an $(r,s)$-type tensor field, then the image of $t$ under the induced connection $D$ is an $(r,s+1)$-type tensor field $Dt$.
Definition 4.
Let $C:u^{i}=u^{i}(t)$ be a parametrized curve on $M$ and
(5.24) $X(t)=x^{i}(t)\left(\frac{\partial}{\partial u^{i}}\right)_{C(t)}.$
$X(t)$ is parallel along $C$ if
(5.25) $\frac{DX}{\operatorname{d}t}=0.$
If the tangent vectors of a curve $C$ are parallel along $C$, then we call $C$ a self-parallel curve, or a geodesic. (5.25) is equivalent with the system of first order differential equations
(5.26) $\frac{\operatorname{d}x^{i}}{\operatorname{d}t}+x^{j}\Gamma^{i}_{jk}\frac{\operatorname{d}u^{k}}{\operatorname{d}t}=0,\quad i=1,\dots,m.$
The tangent vector $X$ at any point of $C$ give rise to a parallel vector field, called the parallel displacement of $X$ along the curve $C$. With (5.17), a geodesic curve $C$ should satisfy the system of second-order differential equations
(5.27) $\frac{\operatorname{d}^{2}u^{i}}{\operatorname{d}t^{2}}+\Gamma^{i}_{jk}\frac{\operatorname{d}u^{j}}{\operatorname{d}t}\frac{\operatorname{d}u^{k}}{\operatorname{d}t}=0,\quad i=1,\dots,m.$
c) Riemannian connections. Following [49, Chapter 5], let us suppose that $M$ is an $m$-dimensional smooth manifold and let $G$ be a symmetric covariant tensor of rank two on $M$. In a local coordinate system $(U,u^{i})$
$G=g_{ij}\operatorname{d}u^{i}\otimes\operatorname{d}u^{j},\quad g_{ij}=g_{ji}.$
We have also
$G(X,Y)=g_{ij}(p)X^{i}Y^{i},\quad X,Y\in T_{p}(M).$
If the local coordinate system $u^{i}$ is changed to $(u^{\prime})^{i}$, then $g_{ij}$ became
(5.28) $g^{\prime}_{ij}=\frac{\partial u^{k}}{\partial{u^{\prime}}^{i}}g_{kl}\frac{\partial u^{l}}{\partial{u^{\prime}}^{j}}.$
If $G$ is a smooth, everywhere nondegenerate symmetric tensor field of rank 2, $M$ is called generalized Riemannian manifold. If $G$ is positive definite, then $M$ is called a Riemannian manifold.
Definition 5.
Suppose $(M,G)$ is an $m$-dimensional generalized Riemannian manifold and $D$ is an affine connection on $M$. If
(5.29) $DG=0,$
then $D$ is called a metric-compatible connection on $(M,G)$.
Condition (5.29) is equivalent with
(5.30) $\operatorname{d}g_{ij}=\omega^{k}_{i}g_{kj}+\omega^{k}_{j}g_{ik},\quad\text{or~{}~{}~{}}\operatorname{d}G=\omega\cdot G+G\cdot\omega^{t}.$
The geometric meaning of metric-compatible connections is that parallel translations preserve the metric [49, page 127].
Let
(5.31) $T^{j}_{ik}:=\Gamma^{j}_{ki}-\Gamma^{j}_{ik}.$
Then
(5.32) $T=T^{j}_{ik}\frac{\partial}{\partial u^{j}}\otimes\operatorname{d}u^{i}\otimes\operatorname{d}u^{k}$
is a $(1,2)$-type tensor, called the torsion tensor of the affine connection $D$ and
$T(X,Y)=D_{X}Y-D_{Y}X-[X,Y].$
If the torsion tensor of the affine connection $D$ is zero, then the connection is said to be a torsion-free connection.
According to [49, page 138]
Theorem 1.
(Fundamental Theorem of Riemannian Geometry) If $M$ is an $m$-dimensional generalized Riemannian manifold, then there exists an unique torsion-free and metric compatible connection on $M$, called the Levi-Civita connection on $M$, or the Riemannian connection of $M$.
Let us denote [49, page 138]:
$\Gamma_{ijk}:=g_{lj}\Gamma^{l}_{ik},\quad w_{ik}:=g_{lk}w^{l}_{i}.$
Also
$\frac{\partial g_{ij}}{\partial u^{k}}=\Gamma_{ijk}+\Gamma_{jik},$
or
$g_{ij,k}:=\partial_{k}g_{ij}-\Gamma^{l}_{ki}g_{lj}-\Gamma^{l}_{kj}g_{il}=0,$
and
(5.33) $\Gamma_{ikj}\!=\!\frac{1}{2}\left(\frac{\partial g_{ik}}{\partial u^{j}}\!+\!\frac{\partial g_{jk}}{\partial u_{i}}-\frac{\partial g_{ij}}{\partial u^{k}}\right),~{}\Gamma^{k}_{ij}=\frac{1}{2}g^{kl}\left(\frac{\partial g_{il}}{\partial u^{j}}\!+\!\frac{\partial g_{jl}}{\partial u^{l}}\!-\!\frac{\partial g_{ij}}{\partial u^{l}}\right),~{}g_{ij}g^{jk}=\delta^{k}_{i}.$
$\Gamma_{ijk}$ ($\Gamma^{i}_{jk}$) is called Christofell’s symbol of first (respectively second) kind. Also $\Omega G$ is an antisymmetric tensor [48, (98)]:
(5.34) $\Omega G+G\Omega^{t}=0.$
Let us introduce [49, pages 14, 142]
$\Omega_{ij}:=\Omega^{k}_{i}g_{kj},$
and the skew symmetric tensor $\Omega_{ij}$ is given by
$\Omega_{ij}=\operatorname{d}\omega_{ij}+\omega^{l}_{i}\wedge\omega_{jl}.$
If
$R_{ijkl}:=\frac{1}{2}R^{h}_{ikl}g_{hj},$
then
$\Omega_{ij}=\frac{1}{2}R_{ijkl}\operatorname{d}u^{k}\wedge\operatorname{d}u^{l},\quad R_{ijkl}=\frac{\partial\Gamma_{ijkl}}{\partial u^{k}}-\frac{\partial\Gamma_{ijk}}{\partial u^{l}}+\Gamma^{h}_{ik}\Gamma_{jhl}-\Gamma^{h}_{il}\Gamma_{jhk}.$
The covariant tensor of rank 4 $R_{ijkl}$ is called the curvature tensor of the generalized Riemannian manifold $M$ and has the properties [49, page 142]:
1. $R_{ijkl}=-R_{jikl}=-R_{ijlk},$
2. $R_{ijkl}+R_{iklj}+R_{iljk}=0,$
3. $R_{ijkl}=R_{klij}.$
5.1.2. Connections on complex manifolds
Several definitions introduced in §5.1.1 for real connections are adapted to the complex case [48]. Essentially, the transpose of $A^{t},~{}A\in M(q,\mathbb{R})$ is replaced by the conjugate transpose (or hermitian transpose) $A^{H}$, $A\in M(q,\mathbb{C})$, also denoted $A^{*}$, $A^{\dagger}$, $A^{+}$.
In Definition 1 $E$ is taken a complex fibre bundle of complex dimension $q$ [48, §5], i.e. the fibres of $E$ are $\mathbb{C}^{q}$ and the structural group is $GL(q,\mathbb{C})$. Then in formula (5.3) the 1-forms $\omega^{\beta}_{\alpha}$ are complex valued and the parallel sections defined in (5.14) are named horizontal lifts.
Using the decomposition $T^{*}_{M}=(T_{M}^{*})^{\prime}+(T_{M}^{*})"$, $D=D^{\prime}+D"$, where $D^{\prime}:{\mathfrak{A}}^{0}(E)\rightarrow{\mathfrak{A}}^{1,0}$ and $D":{\mathfrak{A}}^{0}(E)\rightarrow{\mathfrak{A}}^{0,1}$. The connection $D$ is called compatible with the complex structure if $D"=\bar{\partial}$ [59, page 73].
Instead of Riemannian metric in the real case, a hermitian structure on the complex bundle $E$ is introduced. A hermitian structure on a complex vector space $V$ is a complex-valued function $H(\xi,\eta)$, $\xi,\eta\in V$ such that [48, page 9]
(5.35a) $1$ $~{}~{}H(\lambda_{1}\xi_{1}+\lambda_{2}\xi_{2},\eta)=\lambda_{1}H(\xi_{1},\eta)+\lambda_{2}H(\xi_{2},\eta),\quad\lambda_{1},\lambda_{2}\in\mathbb{C},~{}~{}\xi_{1},\xi_{2},\eta\in V;$
(5.35b) $2$ $~{}~{}\overline{H(\xi,\eta)}=H(\eta,\xi),\quad\xi,\eta\in V.$
$H$ is called positive definite if
$H(\xi,\xi)>0,\quad\xi\not=0.$
A hermitian structure on a complex bundle $\psi:~{}E\rightarrow M$ is a $C^{\infty}$ field of positive definite hermitian structure in the fibers of $E$. A complex vector bundle $\psi:E\rightarrow M$ with a hermitian structure is called hermitian vector bundle. For every frame field $s$ the hermitian structures defines an hermitian matrix
$H_{s}=(H(s_{i},s_{j}))=H_{s}^{H}=\bar{H}_{s}^{t},\quad 1\leq i,j\leq q.$
Under of change of coordinates (5.6), we have instead of (5.28)
$H_{s^{\prime}}=AH_{s}\bar{A}^{t},\quad A\in M(q,\mathbb{C}).$
If condition (5.14) is fulfilled for the hermitian vector bundle, $s$ is called horizontal and condition (5.15) is also fulfilled. The mapping $C\rightarrow\psi^{-1}(C)$ which assign to a point $t\in C$ $s=\lambda^{\alpha}(t)s_{\alpha}$ is called horizontal lifting if (5.15) is satisfied (parallel vector field in Definition 3). Equation (5.7) are satisfied for $A\in M(q,\mathbb{C})$ in (5.6).
Instead of metric compatible connection on generalized Riemannian manifolds, for hermitian vector bundle the connection is called admissible if $H(\xi,\eta)$ remains constant when $\xi,\eta$ are horizontal sections along arbitrary curves. Instead of (5.30), we have
(5.36) $\operatorname{d}G=\omega\cdot H+H\cdot{\omega}^{H},$
where
$H(\xi,\eta)=\sum_{i,k=1}^{n}h_{ik}\xi^{i}\bar{\eta}^{k},\quad\xi=\xi^{i}s_{i},~{}\eta=\eta^{i}s_{i},~{}~{}h_{ik}=H(s_{i},s_{k}).$
The curvature matrix $\Omega$ is introduced as in Definition 5.9 and $\Omega H$ is skew-hermitian, while in the real case is skew-symmetric as in (5.34). A frame field $s$ of a hermitian vector bundle is called unitary if $H_{s}=\mathbb{1}_{q}$ and the connection and curvature matrix are both skew-hermitian.
Now let $M$ be a $m$-dimensional complex manifold and let $\psi:~{}E\rightarrow M$ be a complex vector bundle over $M$ with fiber dimension $q$. If the transition functions $E$ are holomorphic, then $E$ is a holomorphic bundle. If $q=1$ we have a line bundle.
Let $E\rightarrow M$ be a holomorphic vector bundle on the complex manifold $M$ with hermitian metric $h$ defined by the holomorphic reper $f$. The dual bundle $E^{*}\rightarrow M$ has the metric
$h_{E^{*}}(f^{*})=h^{-1}_{E}(f)$
in the dual reper $f^{*}$, and
$\theta_{E^{*}}=-\theta_{E},\quad\Theta_{E^{*}}=-\Theta_{E}.$
Above $\theta_{E}~{}(\Theta_{E})$ are the connection matrix (respectively curvature matrix, denoted in (5.9) with $\Omega$)) of the holomorphic vector bundle $E$ and
(5.37) $\theta_{E}(f)=\partial\log H_{E}(f),\quad\Theta_{E}(f)=\bar{\partial}\theta(f)=\bar{\partial}\partial\log H_{E}(f).$
(5.37) in the case of the projective space give the hermitian metric
$h_{[-1]}(f,f)=(f,f),$
and we find for the tautological line bundle $[-1]$ on ${\mathbb{CP}}^{n}$
$\Theta_{[-1]}(f)=-\frac{(f,f)(\operatorname{d}f,\wedge\operatorname{d}f)-(\operatorname{d}f,f)\wedge(f,\operatorname{d}f)}{(f,f)^{2}},$
We have also the relations
$H_{[-1]}=1+||w||^{2},\quad H_{[1]}=(1+||w||^{2})^{-1}.$
The curvature matrix of the hyperplane line bundle on ${\mathbb{CP}}^{n}$ is
$\Theta_{[1]}=\frac{(1+||w||^{2})\sum\operatorname{d}w_{k}\wedge\operatorname{d}\bar{w}_{k}-\sum\bar{w}_{k}\operatorname{d}w_{k}\wedge\sum w_{k}\operatorname{d}\bar{w}_{k}}{(1+||w||^{2})^{2}}.$
Let $(z^{1}_{U}\dots,z^{m}_{U})$ (respectively $(z^{1}_{V}\dots,z^{m}_{V})$) be local coordinates in $U$ (resp. in $V$). The tangent bundle has as transition functions the Jacobian matrices similar with (5.22)
(5.38) $J_{UV}=\frac{\partial(z^{1}_{U},\dots,z^{m}_{U})}{\partial(z^{1}_{V},\dots,z^{m}_{V})}.$
A section $s\in E$ is holomorphic if its components relative to a chart are holomorphic. A connection such that the connection matrix is a matrix of 1-forms of type (1,0) relative to holomorphic frame field is called a connection of type (1,0). Formulae similar with (5.22), (5.6), (5.23b) hold for 1-forms of type (1,0).
We extract from [29, pages 5, 6] Remark 5 and some considerations.
In the convention $\alpha,\beta,\gamma,\dots$ run from 1 to $n$, while A,B,C,$\dots$ run through $1,\dots,n,$ $\bar{1},\dots,\bar{n}$, [64, p 155], for an almost complex connection without torsion we have the relations
$\Gamma^{\alpha}_{\beta\gamma}=\Gamma^{\alpha}_{\gamma\beta};\quad\bar{\Gamma}^{\alpha}_{\beta\gamma}=\Gamma^{\bar{\alpha}}_{\bar{\beta}\bar{\gamma}}$
and all other $\Gamma^{A}_{BC}$ are zero. For a complex manifold of complex dimension $n$ there are $\frac{n^{2}(n+1)}{2}$ distinct $\Gamma$-s.
If we take into account the hermiticity condition (3.4) in (3.2b) of the metric and the Kählerian restrictions (3.3), the non-zero Christoffel’s symbols $\Gamma$ of the Chern connection (cf. e.g. [5, §3.2], also Levi-Civita connection, cf. e.g. [5, Theorem 4.17]) which appear in (5.33) are determined by the equations, see also e.g. [64, (12) at p 156]
(5.39) $h_{\alpha\bar{\epsilon}}\Gamma^{\alpha}_{\beta\gamma}=\frac{\partial h_{\bar{\epsilon}\beta}}{\partial z_{\gamma}}=\frac{\partial h_{\beta\bar{\epsilon}}}{\partial z_{\gamma}},~{}~{}\alpha,\beta,\gamma,\epsilon=1,\dots,n,m$
and
$\Gamma^{\gamma}_{\alpha\beta}=\bar{h}^{\gamma\bar{\epsilon}}\frac{\partial h_{\beta\bar{\epsilon}}}{\partial z_{\alpha}}=h^{\epsilon\bar{\gamma}}\frac{\partial h_{\beta\bar{\epsilon}}}{\partial z_{\alpha}},\quad\text{where~{} ~{}~{} }h_{\alpha\bar{\epsilon}}h^{\epsilon\bar{\beta}}=\delta_{\alpha\beta}.$
If a hermitian structure $H$ is defined on the holomorphic vector bundle $\psi:~{}E\rightarrow M$, then it has an uniquely defined admissible connection of type (1,0) given by
(5.40) $\omega=\partial H\cdot H^{-1}.$
If $q=1$, then $H=(h),\Omega=(\Omega),h>0$ and
(5.41) $\Omega=-\partial\bar{\partial}\log h.$
$\Omega$ is closed and globally defined.
Chern [48, page 45] calles
(5.42) $\frac{1}{2\pi\operatorname{i}}\Omega$
the curvature form of the connection. The holomorphic line bundle $E\rightarrow M$ is said to be positive if $E$ can be given a metric $h\in\mathbb{C}^{\infty}(M,E^{*}\times\bar{E}^{*})$ such the first Chern class $c_{1}(E)$ is positive.
Remark 5.
Let $M$ be a Kähler  manifold with local complex coordinates $(z^{1},\dots,z^{n})$. Let $\Gamma^{i}_{jk}(z)$ be the holomorphic Christofell’s symbols in the formula of geodesics
(5.43) $\frac{\operatorname{d}^{2}z^{i}}{\operatorname{d}t^{2}}+\Gamma^{i}_{jk}\frac{\operatorname{d}z^{j}}{\operatorname{d}t}\frac{\operatorname{d}z^{k}}{\operatorname{d}t}=0.\quad i=1,\dots,n.$
Let us make in formula (5.43) the change of variables $z^{j}=\xi^{j}+\operatorname{i}\eta^{j}$, $\xi^{i},~{}\eta^{i}\in\mathbb{R}$ and let us introduce the notation $\xi^{j^{\prime}}:=\eta^{j}$, $j^{\prime}:=j+n$, $j=1,\dots,n$.
Then the geodesic equations (5.43) in $(z^{1},\dots,z^{n})\in\mathbb{C}^{n}$ became geodesic equations in the variables $(\xi^{1},\dots,\xi^{n},\xi^{1^{\prime}},\dots,\xi^{n^{\prime}})\in\mathbb{R}^{2n}$
(5.44) $\frac{\operatorname{d}^{2}\xi^{i}}{\operatorname{d}t^{2}}+\tilde{{\Gamma}}^{i}_{jk}\frac{\operatorname{d}\xi^{j}}{\operatorname{d}t}\frac{\operatorname{d}\xi^{k}}{\operatorname{d}t}+2\tilde{{\Gamma}}^{i}_{jk^{\prime}}\frac{\operatorname{d}\xi^{j}}{\operatorname{d}t}\frac{\operatorname{d}\xi^{k^{\prime}}}{\operatorname{d}t}+\tilde{{\Gamma}}^{i}_{j^{\prime}k^{\prime}}\frac{\operatorname{d}\xi^{j^{\prime}}}{\operatorname{d}t}\frac{\operatorname{d}\xi^{k^{\prime}}}{\operatorname{d}t}=0,$
(5.45) $\frac{\operatorname{d}^{2}\xi^{i^{\prime}}}{\operatorname{d}t^{2}}+\tilde{{\Gamma}}^{i^{\prime}}_{jk}\frac{\operatorname{d}\xi^{j}}{\operatorname{d}t}\frac{\operatorname{d}\xi^{k}}{\operatorname{d}t}+2\tilde{{\Gamma}}^{i^{\prime}}_{j^{\prime}k}\frac{\operatorname{d}\xi^{j^{\prime}}}{\operatorname{d}t}\frac{\operatorname{d}\xi^{k}}{\operatorname{d}t}+\tilde{{\Gamma}}^{i^{\prime}}_{j^{\prime}k^{\prime}}\frac{\operatorname{d}\xi^{j^{\prime}}}{\operatorname{d}t}\frac{\operatorname{d}\xi^{k^{\prime}}}{\operatorname{d}t}=0,$
where
(5.46) $\tilde{{\Gamma}}^{i}_{jk}=\tilde{{\Gamma}}^{i^{\prime}}_{j^{\prime}k}=-\tilde{{\Gamma}}^{i}_{j^{\prime}k^{\prime}}=\operatorname{Re}{\Gamma^{i}_{jk}};~{}-\tilde{{\Gamma}}^{i}_{jk^{\prime}}=\tilde{{\Gamma}}^{i^{\prime}}_{jk}=-\tilde{{\Gamma}}^{i^{\prime}}_{j^{\prime}k^{\prime}}=\operatorname{Im}{\Gamma^{i}_{jk}}$
and the real and imaginary parts of $\Gamma^{i}_{jk}$ are functions of $(\xi,\xi^{\prime})\in\mathbb{R}^{2n}$.
We find for the Berry phase (3.11) the expression
(5.47) $\varphi_{B}=\sum_{i,j}(\varphi_{B})^{j}_{i}\!\!=\!-\!\!\sum_{ij}\frac{\operatorname{i}}{2}(\theta^{j}_{i}\!\!-\!\!\bar{\theta}^{j}_{i})\!\!=\!\!\sum_{ij}\operatorname{Im}\Gamma^{j}_{ik}\operatorname{d}\!\xi^{k}\!\!+\!\!\operatorname{Re}\Gamma^{j}_{ik}\operatorname{d}\!\eta^{k}\!\!=\!\!\sum_{ij}\tilde{\Gamma}^{j^{\prime}}_{ik}\operatorname{d}\!\xi^{k}\!\!+\!\!\tilde{\Gamma}^{j}_{ik}\operatorname{d}\!\eta^{k}.$
Proof.
The first part is extracted from [29, Remark 1]. (5.47) is proved with (5.19), (5.40), (5.3). ∎
If $M$ is a complex $m$-dimensional manifold, $M$ is called hermitian if a hermitian structure $H$ is given in its tangent bundle $T(M)$. Then in a natural frame field in local coordinates $z^{1},\dots,z^{m}$
$s_{i}=\frac{\partial}{\partial z^{i}},\quad h_{ik}=H(\frac{\partial}{\partial z^{i}},\frac{\partial}{\partial z^{k}}),\quad H={H}^{H}=(h_{ik}),$
and $H$ is positive definite hermitian. The Kähler  form
$\hat{H}=\frac{\operatorname{i}}{2}\sum h_{ik}\operatorname{d}z^{i}\wedge\operatorname{d}\bar{z}^{k}$
is a real-valued form of type (1,1) and the hermitian manifold $M$ is Kähler  if
$\operatorname{d}\hat{H}=0.$
To a Kähler manifold $(M,\omega_{M})$ it is attached the triple $({L},h,\nabla_{L})$ [81, 37], where ${L}$ is a holomorphic (prequantum) line bundle on $M$, $h$ is a hermitian metric on ${L}$ (taken conjugate linear in the first argument), and $\nabla_{L}$ is a connection compatible with the metric and the Kähler  metric,
(5.48) $\nabla_{L}=\partial+\theta_{L}+\bar{\partial},\quad\theta_{L}=\partial\log\hat{h},$
where $\hat{h}$ is local representative of the hermitian metric $h$; see also [5, Proposition 3.21], where the connection (5.48) is called Chern connection [5, page 31]. With respect to local holomorphic coordinates of the manifold and with respect to a local holomorphic frame for the bundle the metric $h$ can be given as [24, §2]
$h(s_{1},s_{2})(z)=\bar{\hat{h}}(z)\bar{\hat{s}}_{1}(z)\hat{s}_{2}(z),\quad\hat{h}(z)=(e_{z},e_{z})^{-1}.$
where $\hat{s}_{i}$ is a local representing function for the section $s)i$, $i=1,2$, and $\hat{h}$ is a locally defined real-valued function on $M$.
A Kähler  manifold $(M,\omega_{M})$ is quantizable [81] if in local coordinates the curvature of $L$ (5.41) and $\omega_{M}$ are related by the relation
(5.49) $\Omega_{L}=-\operatorname{i}\omega_{M},\quad\text{or}\quad\Omega_{L}=-\partial\bar{\partial}\log{h}.$
Then $\omega_{M}$ is integral, i.e. $\omega_{M}\in H^{2}(M,\mathbb{Z})$ and the first Chen class $c_{1}[{L}]=[\omega_{M}]$ [59, page 141]. $M$ is called a Hodge manifold for compact Kähler  manifolds. Then [][][ ]
$[]:\operatorname{Div}(M)\rightarrow(M,{\mathcal{O}}^{*})$
is a functorial homomorphism between the group of divisors and the Picard group of equivalence classes of $C^{\infty}$ holomorphic line bundles [59, page 133].
Details for quantizable noncompact manifolds can be find in [24, §2, §5], [26, §2.1]
Since $\Omega H$ is of type (1,1), we get
$\Omega H=(\Omega_{ik}),\quad\Omega_{ik}=\sum_{jl}R_{iikj\bar{l}}\sigma^{j}\wedge\bar{\sigma}^{l},$
where $\sigma$ is a base dual with holomorphic base $s$.
Note that in [11], instead of quantization condition (5.49), we used the condition
(5.50) $\Omega_{L}=-2\operatorname{i}\omega_{M}.$
The holomorphic sectional curvature at $(x,\xi)\in(U,T(M))$ is
$R(x,\xi)=2\sum R_{ijkl}\xi^{i}\bar{\xi}^{k}\xi^{j}\bar{\xi}^{l}/(\sum h_{ik}\xi^{i}\bar{\xi}^{k})^{2},$
The (1,1) type form
$\Phi:=\operatorname{Tr}\Omega.$
is called Ricci form.
5.1.3. Holonomy
Following [65], let $\pi:L\rightarrow M$ be a line bundle over $M$ and let ${\mathfrak{L}}={\mathfrak{L}}(M)$ be the set of equivalence classes of line bundles over $M$. ${\mathfrak{L}}$ has a group structure and this group is naturally isomorphic with $H^{2}(M,\mathbb{Z})$. Then there exists locally an unique $\alpha\in{\mathfrak{D}}^{1}$, the connection form, such that [65, (1.4.3)]
(5.51) $\nabla_{\xi}s=2\pi\operatorname{i}<\alpha,\xi>s.$
Comparison of (5.2), (5.5) with (5.51) gives the correspondence $2\pi\operatorname{i}\alpha\rightarrow\omega$ in the notation of connection form in [65] of the and respectively connection matrix in [49].
The construction in (5.51) can be globalized [65, Proposition 1.5.1]. There exists an unique $\Omega\in{\mathfrak{D}}^{2}$ closed such that locally [65, Proposition 1.6.1]
$\operatorname{d}\alpha|_{U}=\Omega|_{U}.$
The closed 2-form $\Omega$ is called the curvature of $(L,\alpha)$ [65, page 104], $\Omega=\operatorname{curv}(L,\alpha)$.
Let $\gamma:[a,b]\rightarrow M$ be a pice-wise smooth curve and there is a linear isomorphism
$P_{\gamma}:L_{\gamma(a)}\rightarrow L_{\gamma(b)}$
called parallel transport along $\gamma$. Then the function $Q:\Gamma\rightarrow\mathbb{C}\setminus\{0\}$ called scalar parallel transport function, where $\Gamma=\Gamma(M)$ is the set of all piece-wise closed curves on $M$
$P_{\gamma}(s)=Q(\gamma)s,\quad~{}~{}\forall s\in L_{p}^{\prime}=L_{p}\setminus\{0\}.$
$Q(\gamma)=\exp(\operatorname{i}\beta)$ is calculated with Stokes’ formula [11]
(5.52) $\beta=\oint_{\gamma}A_{L}=\oint_{\gamma}\operatorname{i}\theta_{L}=\int_{\sigma}\operatorname{d}A_{L},$
where $\sigma$ is a the surface deformation of $\gamma$ [65, page 108] and $\theta_{L}$ is defined in (5.48).
We get [65, Theorem 1.8.1 page 108]
(5.53) $Q(\gamma)=\exp(\operatorname{i}\beta)=\exp(-\oint_{\gamma}\theta_{L})=\exp({\operatorname{i}\int_{\sigma}\omega_{M}}),$
In the convention (5.50) in [11], (5.53) becomes
(5.54) $Q(\gamma)=\exp(\operatorname{i}\beta)=\exp(-\oint_{\gamma}\theta_{L})=\exp(-\oint_{\sigma}\operatorname{d}\theta_{L})=-\oint_{\sigma}\Theta_{L}=\exp({2\operatorname{i}\int_{\sigma}\omega_{M}}),$
With Stokes’ formula (5.52) and (5.53), (5.54) we get in the convention (5.50) [9, (5.4)]
$\operatorname{d}A_{L}=2\omega_{M},$
or
$\operatorname{d}A_{L}=\omega_{M},$
in the convention (5.49).
5.2. Examples
Below for the Heiwsenberg-Weyl group, ${\mathbb{CP}}^{1}$ and ${\mathcal{D}}_{1}$ we follow [27, § 7.2.7].
[…]
Glauber’s coherent states $e_{z}=e^{z{\mbox{\boldmath{$a$}}}^{\dagger}}e_{0}$ have the scalar product
$(e_{\bar{z}},e_{\bar{z}^{\prime}})=e^{z\bar{z}^{\prime}},\quad\omega=\operatorname{i}\operatorname{d}z\wedge\operatorname{d}\bar{z}.$
$\bullet$ ${\mathbb{CP}}^{1}=S^{2}=\text{SU}(2)/\text{U}(1)$. The generators of $\text{SU}(2)$ verify the commutation relations
$[J_{0},J_{\pm}]=\pm J_{\pm};\quad[J_{-},J_{+}]=-2J_{0},$
and the finite dimensional representation of $\text{SU}(2)$ are defined by the action on the extremal weight
[…]
If the CS vectors on $S^{2}$ are introduced as
$e_{z}=e^{z{\mbox{\boldmath{$J$}}}_{+}}e_{j,-j},$
then the scalar product, Kähler  two-form, Berry connection $A_{B}$ and $\operatorname{d}A_{B}$ are respectively
(5.56a) $(e_{\bar{z}},e_{\bar{z}^{\prime}})$ $=(1+z\bar{z}^{\prime})^{2j},\quad\omega_{S^{2}}=2\operatorname{i}j\frac{\operatorname{d}z\wedge\operatorname{d}\bar{z}}{(1+|z|^{2})^{2}},$
(5.56b) $A_{B}$ $=\operatorname{i}j\frac{\bar{z}\operatorname{d}z-z\operatorname{d}\bar{z}}{1+|z|^{2}},\quad\operatorname{d}A_{B}=-2\operatorname{i}j\frac{\operatorname{d}z\wedge\operatorname{d}\bar{z}}{(1+|z|^{2})^{2}}.$
$\bullet$ ${\mathcal{D}}_{1}=\text{SU}(1,1)/\text{U}(1)$. The generators of $\text{SU}(1,1)$ verify the commutation relations
$[K_{0},K_{\pm}]=\pm K_{\pm},\quad[K_{-},K_{+}]=2K_{0}.$
The positive holomorphic discrete series representation corresponds to the action on the extremal weight
[…]
If
$e_{z}=e^{z{\mbox{\boldmath{$K$}}}_{+}}e_{k,k},\quad|z|<1,\quad k=1,\frac{3}{2},2,\frac{5}{2},\dots,$
then the scalar product, Kähler two-form, Berry connection $A_{B}$ and $\operatorname{d}A_{B}$ are respectively
(5.58a) $(e_{\bar{z}},e_{\bar{z^{\prime}}})$ $=(1-z\bar{z}^{\prime})^{-2k},\quad\omega_{{\mathcal{D}}_{1}}=2\operatorname{i}k\frac{\operatorname{d}z\wedge\operatorname{d}\bar{z}}{(1-|z|^{2})^{2}},$
(5.58b) $A_{B}$ $=\operatorname{i}k\frac{\bar{z}\operatorname{d}z-z\operatorname{d}\bar{z}}{1-|z|^{2}},\quad\operatorname{d}A_{B}=-2\operatorname{i}k\frac{\operatorname{d}z\wedge\operatorname{d}\bar{z}}{(1-|z|^{2})^{2}}.$
We remark that (5.58b) has already appeared in (3.56).
Let us denote with $J$ the extremal weight of the representation (5.55) ((5.57)) i.e. $J=-2j$ $(J=2k)$ of $\text{SU}(2)$ (respectively, $\text{SU}(1,1)$). Then we write (5.2) as
(5.59a) $(e_{\bar{z}},e_{\bar{z^{\prime}}})$ $=(1\pm z\bar{z}^{\prime})^{-2J},\quad\omega_{S^{2},{\mathcal{D}}_{1}}=\mp 2\operatorname{i}J\frac{\ \operatorname{d}z\wedge\operatorname{d}\bar{z}}{(1\pm|z|^{2})^{2}},$
(5.59b) $A_{B}$ $=\mp\operatorname{i}J\frac{\bar{z}\operatorname{d}z-z\operatorname{d}\bar{z}}{1\pm|z|^{2}},\quad\operatorname{d}A_{B}=\pm 2\operatorname{i}J\frac{\operatorname{d}z\wedge\ \operatorname{d}\bar{z}}{(1\pm|z|^{2})^{2}},$
where + (-) corresponds to the compact (respectively noncompact) manifold ${\mathbb{CP}}^{1}$ (respectively ${\mathcal{D}}_{1}$).
We also have on $S^{2}$ and ${\mathcal{D}}_{1}$
$\omega_{S^{2},{\mathcal{D}}_{1}}=-\operatorname{d}A_{B}.$
$\bullet$ The Complex Grassmann Manifold $G_{n}(\mathbb{C}^{m+n})=G_{c}/K$ and its noncompact dual $G_{n}/K$, $G_{c}=\text{SU}(n+m)$, $G_{n}=\text{SU}(n,m)$, $K=S(\text{U}(n)\times\text{U}(m))$. See [60, page 452, Type AIII]. Let $Z\in M(n,m,\mathbb{C})$ be Pontrjagin’s coordinates for the compact (noncompact) Grassmann manifold. Then the scalar product of two coherent state vectors is [7, (6.26)]
[…]
for the particular dominant weight [7, §6, Remark 4, (6.25)]
$j=j_{0}=(\underbrace{1,\dots,1}_{n},\underbrace{0,\dots,0}_{m}).$
Formula (5.60) for $\epsilon=1$ for the complex Grassmann manifold appears in [32, (3.20)], [33, (3.6)], using the technique to realize $G_{n}(\mathbb{C}^{m+n})$ as Slater determinat manifold [6].
With formula (3.11) applied to (5.60) and the relation
$\frac{\operatorname{d}}{\operatorname{d}t}\det A=\det A\operatorname{Tr}(A^{-1}\frac{\partial A}{\partial t}),\quad A\in M(n,\mathbb{C}),$
we get
(5.61a) $A_{B}=$ $\frac{\operatorname{i}}{2}\operatorname{Tr}[(\operatorname{d}ZZ^{+}-Z\operatorname{d}Z^{+})(\mathbb{1}_{n}+\epsilon ZZ^{+})^{-1}],$
(5.61b) $\operatorname{d}A_{B}=$ $-\operatorname{i}\operatorname{Tr}[\operatorname{d}Z(\mathbb{1}_{m}+\epsilon Z^{+}Z)^{-1}\wedge\operatorname{d}Z^{+}(\mathbb{1}_{n}+\epsilon ZZ^{+})^{-1}].$
Formula (5.61a) of the Berry connection on the complex Grassmann manifold $G_{n}(\mathbb{C}^{n+m})$ corresponding to the scalar product (1) was obtained in [35, (5.17)]
(5.62) $A_{B}=\frac{\operatorname{i}}{2}\operatorname{Tr}[(Z^{+}\operatorname{d}Z-\operatorname{d}Z^{+}Z)(\mathbb{1}_{m}+Z^{+}Z)^{-1}],$
where $Z\in M(m,n,\mathbb{C})$, which is identical with (5.60) if $m\leftrightarrow n$.
Now we apply formula (3.2b) for $G_{c}/K$ and $G_{n}/K$ for the scalar product (5.60) and we get
(5.63) $\omega=\operatorname{i}\operatorname{Tr}[(\mathbb{1}_{n}+\epsilon ZZ^{+})^{-1}\operatorname{d}Z\wedge(\mathbb{1}_{m}+\epsilon Z^{+}Z)^{-1}\operatorname{d}Z^{+}].$
Comparing (5.61b) and (5.63), it follows that on $G_{c}/K$ and $G_{n}/K$ we have the relation
(5.64) $\operatorname{d}A_{B}=-\omega.$
Formula (5.61b) for $\epsilon=1$ on $G_{n}(\mathbb{C}^{n+m})$ appears in [35, page 1005], where it was emphasized that it is the explicit realization of the two-form $V$ of Simon [83].
We recall that the invariant metric on $X_{c}$ ($X_{n}$) [7, (6.10)] is
$\operatorname{d}s^{2}=k\operatorname{Tr}[\operatorname{d}Z(\mathbb{1}_{m}+\epsilon Z^{+}Z)^{-1}\operatorname{d}Z^{+}(\mathbb{1}_{n}+\epsilon ZZ^{+})^{-1}],\quad Z\in M(n,m,\mathbb{C}).$
The equation of geodesics on $X_{c,n}$ [7, (6.13)]
$\frac{\operatorname{d}^{2}}{\operatorname{d}t^{2}}-2\epsilon\frac{\operatorname{d}Z}{\operatorname{d}t}Z^{+}(\mathbb{1}_{n}+\epsilon ZZ^{+})^{-1}\frac{\operatorname{d}Z}{\operatorname{d}t}=0$
has the solution $Z=Z(tB)$
$Z=Z(B)=B\frac{\text{ta}\sqrt{B^{+}B}}{\sqrt{B^{+}B}},\quad B\in M(n,m,\mathbb{C}),$
with the initial condition $Z(0)=B$, where $\text{ta}=\tan$ ($\tanh$) for $X_{c}$ (respectively, $X_{n}$) .
Equations (5.61a) ((5.61b)) for $\frac{\text{SU}(m+n)}{\text{SU}(m)\times\text{SU}(n)}$ (respectively $\frac{\text{SU}(m,n)}{\text{SU}(m)\times\text{SU}(n)}$) of $\varphi_{B}=\oint A_{B}$, $\operatorname{d}A_{B}$ have been obtained in [31, (62), (63)].
In [11] in the relation $A_{L}=\operatorname{i}\theta_{L}$ the connection matrix $\theta_{L}$ corresponds to the Hermitian metric on the dual of the tautological line bundle on the Grassmann manifold $\hat{h}_{L}(Z)=\det(\mathbb{1}_{n}+ZZ^{+})^{-1}$. The Berry connection which corresponds to $A_{L}$ is [11, (5.2)]
(5.65) $A_{B}=\frac{\operatorname{i}}{2}\operatorname{Tr}[(\operatorname{d}ZZ^{+}-Z\operatorname{d}Z^{+})(\mathbb{1}_{n}+ZZ^{+})^{-1}].$
The corresponding two-form on $G_{n}(\mathbb{C}^{n+m})$ is [11, (5.3)]
$\omega=\frac{\operatorname{i}}{2}\operatorname{Tr}[\operatorname{d}Z(\mathbb{1}_{m}+Z^{+}Z)^{-1}\wedge\operatorname{d}Z^{+}(\mathbb{1}_{n}+ZZ^{+})^{-1}].$
We reported [9, (5.5)] the holomorphic connection on the compact complex Grassmann manifold and its non-compact dual
$A_{L}=\operatorname{i}\operatorname{Tr}[\operatorname{d}ZZ^{+}(\mathbb{1}_{n}+\epsilon ZZ^{+})^{-1}]$
corresponding to the hyperplan line bundle, the dual of the tautological line bundle $L$ on the Grassmannian and its non-compact dual [9, (5.6)]
(5.66) $\hat{h}_{L}=\det(\mathbb{1}_{n}+\epsilon ZZ^{+})^{-\epsilon},$
the Kähler  two-form is [9, (5.7)], where $\mathbb{1}_{n}$ should be replaced with $\mathbb{1}_{m}$
$\omega=\frac{\operatorname{i}}{2}\operatorname{Tr}[\operatorname{d}Z(\mathbb{1}_{m}+\epsilon Z^{+}Z)^{-1}\wedge\operatorname{d}Z^{+}(\mathbb{1}_{n}+\epsilon ZZ^{+})^{-1}],$
while the Berry connection is [9, (5.8)]
(5.67) $A_{B}=\frac{\operatorname{i}}{2}\operatorname{Tr}[(\operatorname{d}ZZ^{+}-Z\operatorname{d}Z^{+})(\mathbb{1}_{n}+\epsilon ZZ^{+})^{-1}].$
$\bullet$ The complex projective space ${\mathbb{CP}}^{n}$ and ${\mathbb{CP}}^{n,1}$. We recall that the ray space is defined as
${\mathbb{CP}}^{n}={\mathbb{P}}(\mathbb{C}^{n})=S^{2n+1}/\approx=\frac{\text{SU}(n+1)}{S(\text{U}(n)\times\text{U}(1))},\quad\text{where~{}~{}~{}}x\approx y\leftrightarrow x=\lambda y,\lambda\in\text{U}(1).$
Also we have
${\mathbb{CP}}^{n}=\mathbb{C}^{n+1}/\approx\quad\text{where~{}~{}~{}}x\approx y\leftrightarrow x=\lambda y,\lambda\in\mathbb{C}^{*}=\mathbb{C}\setminus 0.$
Also we have the dual space
${\mathbb{CP}}^{n,1}=\frac{\text{SU}(n,1)}{S(\text{U}(n)\times\text{U}(1))}.$
In [33, (24)], [34, (5.8)] we have proved that for ${\mathbb{CP}}^{N}$ and $M(1,N)\ni Z=(Z_{1},\dots,Z_{N})$ that
(5.68c) $(\mathbb{1}_{N}+Z^{+}Z)^{-1}_{\alpha,\beta}$ =$(\mathbb{1}_{N}+|Z|^{2})^{-1}\left\{\begin{array}[]{l}1+|Z|^{2}-|Z_{\alpha}|^{2},\quad\alpha=\beta\\ -\bar{Z}_{\alpha}Z_{\beta},\quad\quad\quad\quad\quad\alpha\not=\beta\end{array}\right.$
(5.68d) $=\frac{(1+|Z|^{2})\delta_{\alpha\beta}-\bar{Z}_{\alpha}Z_{\beta}}{1+|Z|^{2}},$
where
$|Z|^{2}=|Z_{1}|^{2}+|Z_{2}|^{2}+\cdots|Z_{N}|^{2}.$
We recall that the Fubini-Studdy metric on ${\mathbb{CP}}^{N}$ is
(5.69) $g_{\alpha\beta}=\frac{(1+|Z|^{2})\delta_{\alpha\beta}-\bar{Z}_{\alpha}Z_{\beta}}{(1+|Z|^{2})^{2}}.$
For ${\mathbb{CP}}^{N,1}$ we have (5.2) replacing (5.2) for ${\mathbb{CP}}^{N}$ and we can write together ${\mathbb{CP}}^{N}$ ($\epsilon=1$) and ${\mathbb{CP}}^{N,1}$ ($\epsilon=-1$) as
(5.70c) $(\mathbb{1}_{N}+\epsilon Z^{+}Z)^{-1}_{\alpha,\beta}$ =$(\mathbb{1}_{N}+\epsilon|Z|^{2})^{-1}\left\{\begin{array}[]{l}(1+\epsilon|Z|^{2}),\quad\alpha=\beta\\ (-\epsilon\bar{Z}_{\alpha}Z_{\beta}),\quad\quad\quad\quad\quad\alpha\not=\beta\end{array}\right.$
(5.70d) $=\frac{(1+\epsilon|Z|^{2})\delta_{\alpha\beta}-\epsilon\bar{Z}_{\alpha}Z_{\beta}}{1+\epsilon|Z|^{2}}.$
Now we introduce (5.2) into (5.63) and we get
(5.71) $\omega_{{\mathbb{CP}}^{n},{\mathbb{CP}}^{n,1}}(Z,\bar{Z})=\operatorname{i}g_{\alpha\beta}\operatorname{d}Z_{\alpha}\wedge\operatorname{d}\bar{Z}_{\beta}=\operatorname{i}\frac{(1+\epsilon|Z|^{2})\delta_{\alpha\beta}-\epsilon\bar{Z}_{\alpha}{Z}_{\beta}}{(1+\epsilon`|Z|^{2})^{2}}\operatorname{d}Z_{\alpha}\wedge\operatorname{d}\bar{Z}_{\beta},$
which is the Fubini-Study Kähler  two - form (5.69) on ${\mathbb{CP}}^{n}$ (respectively, its non-compact dual ${\mathbb{CP}}^{n,1}$, sometimes called the hyperbolic space and denoted ${\mathbb{H}}^{n}$ [87, page 67] ). The condition on $Z$ for $\epsilon=-1$, $(n,m)=(1,n)$, is [31, (23)]
$01-|Z|^{2}>$ 0 .
If equation (5.71) we put $n=1$ we regain (5.56a) for $\epsilon=1$ ( $\epsilon=-1$) $S^{2}$, (5.56a), $j=\frac{1}{2}$ (${\mathcal{D}}_{1}$,(5.58a)), respectively, $k=\frac{1}{2}$).
We recall also the definition of the tautological line bundle $[-1]$
[…]
associated to the transition functions
$g_{ij}([z])=\{\frac{z_{i}}{z_{j}}\},\quad[z]\in U_{i}\cap U_{j}.$
$[1]$ is the hyperplane bundle, the dual of the tautological bundle $[-1]$.
The tautological line bundle does not have global holomorphic sections not identically 0.
Formula (5.61a) particularised for ${\mathbb{CP}}^{n}$ and ${\mathbb{CP}}^{n,1}$ reads
(5.73) $A_{B}=\frac{\operatorname{i}}{2}\frac{\bar{z}_{i}\operatorname{d}z_{i}-z_{i}\operatorname{d}\bar{z}_{i}}{1+\epsilon|z|^{2}},$
which for $\epsilon=1$ is [77, (10)]. If we differentiate (5.73), we got
(5.74) $\operatorname{d}A_{B}=\operatorname{i}\frac{\epsilon\bar{z}_{i}z_{j}-(1+\epsilon|z|^{2})\delta_{ij}}{(1+\epsilon|z|^{2})^{2}}\operatorname{d}z_{i}\wedge\operatorname{d}\bar{z}_{j},$
which is (5.61b) particularised for $m=1$ with (5.2).
Applying (5.41) to (5.66) for ${\mathbb{CP}}^{n},{\mathbb{CP}}^{n,1}$, we get
$\Omega_{{\mathbb{CP}}^{n},{\mathbb{CP}}^{n,1}}=\frac{-\epsilon\bar{Z}_{\alpha}Z_{\beta}+(1+\epsilon|Z|^{2})\delta_{\alpha\beta}}{(1+\epsilon\mathbb{Z}|^{2})}\operatorname{d}Z_{\alpha}\wedge\operatorname{d}\bar{Z}_{\beta}=-\operatorname{i}\omega_{{\mathbb{CP}}^{n},{\mathbb{CP}}^{n,1}},$
which is the quantizablity condition (5.49) of ${\mathbb{CP}}^{n}$ (respectively ${\mathbb{CP}}^{n,1}$).
Acknowledgements
This research was conducted in the framework of the ANCS project program PN 19 06 01 01/2019.